### v1

# ON EQUICONTINUITY OF MAPPINGS WITH VARIABLE DEFINITION DOMAINS

Abstract
The paper is devoted to the study of the global behavior of mappings. We consider families mappings with different definition domains which satisfy moduli inequalities of Poletskii type. Under some additional assumptions we have proved that such families are uniformly equicontinuous.
2010 Mathematics Subject Classification: Primary 30C65; Secondary 31A15, 31B25
1 Introduction
This paper is devoted to the study of mappings satisfying upper distortion estimates of the modulus of families of paths. We consider the case when definition domains of mappings may be different. The main subject of this study is distortion estimates for the distance under mappings. For fixed domains, this problem has been repeatedly studied by many authors, including both homeomorphisms and mappings with a branching. In particular, we could cite papers on this matter, see, e.g., [Cr1][Cr2], [MRSY], [RSY1][RSY2], [RS1], [Sev1] and [SevSkv]. We note that we draw heavily on the ideas of Suvorov [Suv], who considered sequences of domains and the theory of prime ends for them. At the same time, the present paper is not, generally speaking, concerned with the theory of prime ends.
Let us recall some definitions. A Borel function $\rho:{\mathbb{R}}^{n}\,\rightarrow[0,\infty]$ is called admissible for the family $\Gamma$ of paths $\gamma$ in ${\mathbb{R}}^{n},$ if the relation
$\int\limits_{\gamma}\rho(x)\,|dx|\geqslant 1$
holds for all (locally rectifiable) paths $\gamma\in\Gamma.$ In this case, we write: $\rho\in{\rm adm}\,\Gamma.$ Let $p\geqslant 1,$ then $p$-modulus of $\Gamma$ is defined by the equality
$M_{p}(\Gamma)=\inf\limits_{\rho\in\,{\rm adm}\,\Gamma}\int\limits_{{\mathbb{R}}^{n}}\rho^{p}(x)\,dm(x)\,.$
We set $M(\Gamma):=M_{n}(\Gamma).$ Let $x_{0}\in{\mathbb{R}}^{n},$ $0<r_{1}<r_{2}<\infty,$
$S(x_{0},r)=\{x\,\in\,{\mathbb{R}}^{n}:|x-x_{0}|=r\}\,,\quad B(x_{0},r)=\{x\,\in\,{\mathbb{R}}^{n}:|x-x_{0}|<r\}$ (1.1)
and
$A=A(x_{0},r_{1},r_{2})=\left\{x\,\in\,{\mathbb{R}}^{n}:r_{1}<|x-x_{0}|<r_{2}\right\}\,.$ (1.2)
Given sets $E,$ $F\subset\overline{{\mathbb{R}}^{n}}$ and a domain $D\subset{\mathbb{R}}^{n}$ we denote by $\Gamma(E,F,D)$ a family of all paths $\gamma:[a,b]\rightarrow\overline{{\mathbb{R}}^{n}}$ such that $\gamma(a)\in E,\gamma(b)\in\,F$ and $\gamma(t)\in D$ for $t\in(a,b).$ Let $S_{i}=S(x_{0},r_{i}),$ $i=1,2,$ where spheres $S(x_{0},r_{i})$ centered at $x_{0}$ of the radius $r_{i}$ are defined in (1.1). Given $p\geqslant 1,$ a mapping $f:D\rightarrow\overline{{\mathbb{R}}^{n}}$ is called a ring $Q$-mapping at a point $x_{0}\in{\mathbb{R}}^{n}$ with respect to $p$-modulus, if the condition
$M_{p}(f(\Gamma(S_{1},S_{2},D)))\leqslant\int\limits_{A\cap D}Q(x)\cdot\eta^{p}(|x-x_{0}|)\,dm(x)$ (1.3)
holds for all $0<r_{1}<r_{2}<d_{0}:=\sup\limits_{x\in D}|x-x_{0}|$ and any Lebesgue measurable function $\eta:(r_{1},r_{2})\rightarrow[0,\infty]$ such that
$\int\limits_{r_{1}}^{r_{2}}\eta(r)\,dr\geqslant 1\,.$ (1.4)
The inequalities of the form (1.3) were established for many well-known classes of mappings, see e.g. [MRV, Theorem 7.1], [Va, Definition 13.1]), [MRSY, Theorems 8.1, 8.5] and [Sal, Corollary 2].
Recall that a mapping $f:D\rightarrow{\mathbb{R}}^{n}$ is called discrete if the pre-image $\{f^{-1}\left(y\right)\}$ of each point $y\,\in\,{\mathbb{R}}^{n}$ consists of isolated points, and is open if the image of any open set $U\subset D$ is an open set in ${\mathbb{R}}^{n}.$ A mapping $f:D\rightarrow{\mathbb{R}}^{n}$ is called closed if $f(A)$ is closed in $f(D)$ whenever $A$ is closed in $D.$ Later, in the extended space $\overline{{{\mathbb{R}}}^{n}}={{\mathbb{R}}}^{n}\cup\{\infty\}$ we use the spherical (chordal) metric $h(x,y)=|\pi(x)-\pi(y)|,$ where $\pi$ is a stereographic projection $\overline{{{\mathbb{R}}}^{n}}$ onto the sphere $S^{n}(\frac{1}{2}e_{n+1},\frac{1}{2})$ in ${{\mathbb{R}}}^{n+1},$ namely,
$h(x,\infty)=\frac{1}{\sqrt{1+{|x|}^{2}}}\,,\quad h(x,y)=\frac{|x-y|}{\sqrt{1+{|x|}^{2}}\sqrt{1+{|y|}^{2}}}\,,\quad x\neq\infty\neq y$
(see [Va, Definition 12.1]). In what follows, ${\rm Int\,}A$ denotes the set of inner points of the set $A\subset\overline{{\mathbb{R}}^{n}}.$ Recall that the set $U\subset\overline{{\mathbb{R}}^{n}}$ is neighborhood of the point $z_{0},$ if $z_{0}\in{\rm Int\,}A.$
Given $E_{1},E_{2}\subset\overline{{\mathbb{R}}^{n}}$ we set
$h(E_{1},E_{2})=\inf\limits_{x\in E_{1},y\in E_{2}}h(x,y)\,.$
Let $I$ be a fixed set of indices and let $D_{i},$ $i\in I,$ be some number of domains. Following [NP, Sect. 2.4], we say that a family of domains $\{D_{i}\}_{i\in I}$ is equi-uniform with respect to $p$-modulus if for any $r>0$ there exists a number $\delta>0$ such that the inequality
$M_{p}(\Gamma(F^{\,*},F,D_{i}))\geqslant\delta$ (1.5)
holds for any $i\in I$ and any continua $F,F^{*}\subset D_{i}$ such that $h(F)\geqslant r$ and $h(F^{\,*})\geqslant r.$ Set
$q_{x_{0}}(r):=\frac{1}{\omega_{n-1}r^{n-1}}\int\limits_{|x-x_{0}|=r}Q(x)\,dS\,$,
where $dS$ is the element of the surface area of $S,$ and
$q^{\,\prime}_{x_{0}}(r):=\frac{1}{\omega_{n-1}r^{n-1}}\int\limits_{|x-x_{0}|=r}Q^{\,\prime}(x)\,dS\,$,
$Q^{\,\prime}(x)=\max\{Q(x),1\}.$ A domain $D\subset{\mathbb{R}}^{n}$ is called locally connected at the point $x_{0}\in\partial D,$ if for any neighborhood $U$ of point $x_{0}$ there is a neighborhood $V\subset U$ of the same point such that $V\cap D$ is connected. A domain $D$ is called locally connected on $\partial D,$ if this domain is such at each point of its boundary.
Let $D_{m},$ $m=1,2,\ldots,$ be a sequence of domains in ${\mathbb{R}}^{n},$ containing a fixed point $A_{0}.$ If there exists a ball $B(A_{0},\rho),$ $\rho>0,$ belonging to all $D_{m},$ then the kernel of the sequence $D_{m},$ $m=1,2,\ldots,$ with respect to $A_{0}$ is the largest domain $D_{0}$ containing $x_{0}$ and such that for each compact set $E$ belonging to $D_{0}$ there is $N>0$ such that $E$ belongs to $D_{m}$ for all $m\geqslant N.$ A largest domain is one which contains any other domain having the same property. A sequence of domains $D_{m},$ $m=1,2,\ldots,$ converges to a kernel $D_{0}$ if any subsequence of $D_{m}$ has $D_{0}$ as its kernel.
Let $D_{m},$ $m=1,2,\ldots,$ be a sequence of domains which converges to a kernel $D_{0}.$ Then $D_{m}$ will be called regular with respect to $D_{0},$ if $D_{m}\subset D_{0}$ for all $m\in{\mathbb{N}}$ and, for every $x_{0}\in\overline{D_{0}}$ and for every neighborhood $U$ of $x_{0}$ there is a neighborhood $V$ of $x_{0},$ $V\subset U,$ and $M=M(U,x_{0})\in{\mathbb{N}}$ such that $V\cap D_{m}$ is a non-empty connected set for every $m\geqslant M(U,x_{0}).$
Given $p\geqslant 1,$ a number $\delta>0,$ a fixed domain $D_{0}\subset{\mathbb{R}}^{n},$ $n\geqslant 2,$ a sequence of domains $\mathfrak{D}=\{D_{m}\}_{m=1}^{\infty}$ the kernel of which is $D_{0},$ a non-degenerate continuum $A\subset D_{0},$ $A\subset\bigcap\limits_{m=1}^{\infty}D_{m},$ and a Lebesgue measurable function $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$, we denote by $\mathfrak{F}_{Q,A,p,\delta}(D_{0},\mathfrak{D})$ some a family of ring $Q$-homeomorphisms $f_{m}:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}$ satisfying (1.3)–(1.4) at any point $x_{0}\in\overline{D_{0}}$ such that $h(f_{m}(A))\geqslant\delta$ and $h(\overline{{\mathbb{R}}^{n}}\setminus f_{m}(D_{m}))\geqslant\delta.$
Theorem 1.1.   Let $p\in(n-1,n]$ and let $f_{m}\in\mathfrak{F}_{Q,A,p,\delta}(D_{0},\mathfrak{D}),$ $m=1,2,\ldots,$ be a sequence such that:
1) the sequence of domains $D_{m}$ is regular with respect to $D_{0};$
2) for every $m\in{\mathbb{N}},$ a domain $D_{m}$ is locally connected on its boundary;
3) the family $f_{m}(D_{m})$ is equi-uniform with respect to $p$-modulus over all $m\in{\mathbb{N}};$
4) at least one of two following conditions hold: $Q$ has a finite mean oscillation in $\overline{D_{0}},$ or
$\int\limits_{0}^{\beta(x_{0})}\frac{dt}{t^{\frac{n-1}{p-1}}q_{x_{0}}^{\,\prime\,\frac{1}{p-1}}(t)}=\infty$ (1.6)
for every $x_{0}\in\overline{D_{0}}$ and some $\beta(x_{0})>0.$
Then the family $f_{m},$ $m=1,2,\ldots,$ is uniformly equicontinuous in $\mathfrak{D},$ i.e., for any $\varepsilon>0$ there is $\delta=\delta(\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(y))<\varepsilon$ whenever $x,y\in D_{m},$ $|x-y|<\delta$ and $m\in{\mathbb{N}}.$
Moreover, there is a subsequence $f_{m_{k}},$ $k=1,2,\dots,$ which converges to $f$ locally uniformly in $D_{0}.$ In this case, $f$ has a continuous boundary extension $f:\overline{D_{0}}\rightarrow\overline{{\mathbb{R}}^{n}}.$
In addition, assume that $p=n$ and:
5) there is $r>0$ such that $h(E)\geqslant r$ whenever $E$ is any component of $\partial f_{m}(D_{m});$
6) for each point $x_{0}\in D_{0}$ and for every $0<r_{1}<r_{2}<r_{0}:=\sup\limits_{x\in D_{0}}|x-x_{0}|$ there is a set $E_{1}\subset[r_{1},r_{2}]$ of a positive linear Lebesgue measure such that the function $Q$ is integrable with respect to $\mathcal{H}^{n-1}$ over the spheres $S(x_{0},r)$ for every $r\in E_{1}.$ (In particular, the latter holds if $Q$ is a locally integrable function).
Then $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots$ and some $\delta_{1}>0.$
We also prove a result similar to Theorem 1 for non-injective mappings.
Given $p\geqslant 1,$ a domain $D_{0}\subset{\mathbb{R}}^{n},$ a set $E\subset\overline{{\mathbb{R}}^{n}}$ and a number $\delta>0$, a fixed domain $D_{0}\subset{\mathbb{R}}^{n},$ $n\geqslant 2,$ a sequence of domains $\mathfrak{D}=\{D_{m}\}_{m=1}^{\infty}$ the kernel of which is $D_{0}$ and a Lebesgue measurable function $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$ we denote by $\mathfrak{R}_{Q,\delta,p,E}(D_{0},\mathfrak{D})$ the family of all open discrete closed ring mappings $f:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}\setminus E$ satisfying (1.3)–(1.4) at any point $x_{0}\in\overline{D_{0}}$ such that the following condition holds: for any domain $D^{\,\prime}_{m}:=f_{m}(D_{m})$ there exists a continuum $K_{m}\subset D^{\,\prime}_{m}$ such that $h(K_{m})\geqslant\delta$ and $h(f_{m}^{\,-1}(K_{m}),\partial D_{m})\geqslant\delta>0.$ The following statement holds.
Theorem 1.2.Let $p\in(n-1,n]$ and let $f_{m}\in\mathfrak{R}_{Q,\delta,p,E}(D_{0},\mathfrak{D}),$ $m=1,2,\ldots,$ be a sequence such that the conditions 1)–4) from Theorem 1 hold. Let $E$ be a set of a positive capacity for $p=n,$ and $E$ is arbitrary closed set for n1<p<n.n-1<p<n.
Then the family $f_{m},$ $m=1,2,\ldots,$ is uniformly equicontinuous in $\mathfrak{D},$ i.e., for any $\varepsilon>0$ there is $\delta=\delta(\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(y))<\varepsilon$ whenever $x,y\in D_{m},$ $|x-y|<\delta$ and $m\in{\mathbb{N}}.$
Moreover, there is a subsequence $f_{m_{k}},$ $k=1,2,\dots,$ which converges to $f$ locally uniformly in $D_{0}.$ In this case, $f$ has a continuous boundary extension $f:\overline{D_{0}}\rightarrow\overline{{\mathbb{R}}^{n}}.$ Besides that, if $p=n$ and conditions 5)–6) of Theorem 1 hold, then for any continuum $A\subset D_{0},$ $A\subset\bigcap\limits_{m=1}^{\infty}D_{m},$ such that $h(f_{m}(A))\geqslant\delta$ for any $m=1,2,\ldots,$ there exists $\delta_{1}>0$ such that $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots.$
2 Preliminaries
The following statement is true, see [Sev2, Lemma 2.1].
Lemma 2.1.  Let $p\geqslant 1,$ $f:D\rightarrow\overline{{\mathbb{R}}^{n}}$ be an open discrete ring $Q$-mapping at the point $b\in\partial D$ with respect to the $p$-modulus, $b\neq\infty,$ $f(D)=D^{\,\prime}$, a domain $D$ is locally connected at the point $b,$ $C(f,\partial D)\subset\partial D^{\,\prime}$ and the domain $D^{\,\prime}$ is strongly accessible with respect to the $p$-modulus at least at one point $y\in C(f,b),$ where $C(f,b)$ is a cluster set of $f$ at $b.$ Suppose that there is $\varepsilon_{0}>0$ and some positive measurable function $\psi:(0,\varepsilon_{0})\rightarrow(0,\infty)$ such that
$0<I(\varepsilon,\varepsilon_{0})=\int\limits_{\varepsilon}^{\varepsilon_{0}}\psi(t)\,dt<\infty$ (2.1)
for any $\varepsilon\in(0,\varepsilon_{0})$ and, in addition,
$\int\limits_{A(\varepsilon,\varepsilon_{0},b)}Q(x)\cdot\psi^{\,p}(|x-b|)\ dm(x)=o(I^{p}(\varepsilon,\varepsilon_{0}))\,$, (2.2)
where $A:=A(b,\varepsilon,\varepsilon_{0})$ is defined in (1.2). Then $C(f,b)=\{y\}.$
Observe that, $C(f,\partial D)\subset\partial D^{\,\prime}$ for homeomorphisms of $D$ onto $D^{\,\prime}$, see e.g. [MRSY, Proposition 13.5].
We will prove the main assertions of the paper in a somewhat more general form. In this regard, consider the following definition. We will say that a point $x_{0}$ is a limit point of a family of domains $\mathfrak{D}$ in ${\mathbb{R}}^{n}$ if there exists a sequence $D_{m},$ $m=1,2,\ldots,$ in $\mathfrak{D}$ and a sequence $x_{m}\in D_{m},$ such that $x_{m}\rightarrow x_{0}$ as $m\rightarrow\infty.$ Let $\overline{\mathfrak{D}}$ be a set of all such points $x_{0}.$
Given $p\geqslant 1,$ a number $\delta>0,$ a family $\mathfrak{D}$ of domains in ${\mathbb{R}}^{n},$ $n\geqslant 2,$ and a Lebesgue measurable function $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$, we denote by $\mathfrak{H}^{p}_{Q,\delta}(\mathfrak{D})$ a family of all homeomorphisms $f:D\rightarrow\overline{{\mathbb{R}}^{n}}$ defined in some $D\in\mathfrak{D}$ and satisfying (1.3)–(1.4) at any point $x_{0}\in\overline{\mathfrak{D}}$ for which there exists a continuum $A=A_{f}\subset D$ with $h(f(A))\geqslant\delta.$
The following lemma asserts the equicontinuity of a certain family of mappings at a fixed point. Unlike the main result of Theorem 1, this assertion does not assume the convergence of the sequence of domains to its kernel, and the corresponding domains in which the mappings are defined can be arbitrary. The following statement is true.
Lemma 2.2.   Let $p\in(n-1,n],$ let $x_{0}\in\overline{\mathfrak{D}}$ and let $f_{m}\in\mathfrak{H}^{p}_{Q,\delta}(\mathfrak{D}),$ $m=1,2,\ldots,$ $f_{m}:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}$ be a sequence such that:
1) for every $m\in{\mathbb{N}},$ $D_{m}$ is locally connected at $\partial D_{m};$
2) for every neighborhood $U$ of $x_{0}$ there is a neighborhood $V$ of $x_{0},$ $V\subset U,$ and $M=M(U,x_{0})\in{\mathbb{N}}$ such that $V\cap D_{m}$ is connected for every $m\geqslant M(U,x_{0});$
3) there is $\varepsilon_{0}>0$ such that $A_{m}:=A_{f_{m}}\subset{\mathbb{R}}^{n}\setminus B(x_{0},\varepsilon_{0}),$ where $A_{f_{m}}$ is a continuum corresponding to the definition of the class $\mathfrak{H}^{p}_{Q,\delta}(\mathfrak{D})$ for $f_{m};$
4) the family $f_{m}(D_{m})$ is equi-uniform with respect to $p$-modulus over all $m\in{\mathbb{N}};$
5) there exists $\varepsilon_{0}=\varepsilon_{0}(x_{0})>0$ and a Lebesgue measurable function $\psi:(0,\varepsilon_{0})\rightarrow[0,\infty]$ such that
$I(\varepsilon,\varepsilon_{0}):=\int\limits_{\varepsilon}^{\varepsilon_{0}}\psi(t)\,dt<\infty\quad\forall\,\,\varepsilon\in(0,\varepsilon_{0})\,,\quad I(\varepsilon,\varepsilon_{0})\rightarrow\infty\quad\text{as}\quad\varepsilon\rightarrow 0\,$, (2.3)
while
$\int\limits_{A(x_{0},\varepsilon,\varepsilon_{0})}Q(x)\cdot\psi^{\,p}(|x-x_{0}|)\,dm(x)=o(I^{p}(\varepsilon,\varepsilon_{0}))\,$, (2.4)
as $\varepsilon\rightarrow 0,$ where $A(x_{0},\varepsilon,\varepsilon_{0})$ is defined in (1.2).
Then the family $f_{m},$ $m=1,2,\ldots,$ is equicontinuous with respect to $x_{0},$ i.e., for any $\varepsilon>0$ there is $\delta=\delta(x_{0},\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(x^{\,\prime}))<\varepsilon$ whenever $x,x^{\,\prime}\in B(x_{0},\delta)\cap D_{m}$ and $m\in{\mathbb{N}}.$
Proof.   We will mainly use the methodology that has been used for mappings of a fixed domain, see, e.g., [SevSkv]. Assume the contrary. Then there exists $a>0$ such that for $\delta=1/k,$ $k=1,2,\ldots,$ there are $m_{k}\in{\mathbb{N}}$ and $x_{k},x^{\,\prime}_{k}\in B(x_{0},1/k)\cap D_{m_{k}}$ such that $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a.$ We may consider that the sequence $m_{k}$ is increasing by $k.$ Otherwise, $m_{k}=k_{0}$ for sufficiently large $k\in{\mathbb{N}}$ and some $k_{0}\in{\mathbb{N}},$ besides that $x_{0}\in\overline{D_{k_{0}}}$ because $f_{k_{0}}$ is defined at $x_{k}\cap D_{k_{0}}$ and $x_{k}\rightarrow x_{0}$ as $k\rightarrow\infty.$ Now, from the assumption above we have that $h(f_{k_{0}}(x_{k}),f_{k_{0}}(x^{\,\prime}_{k}))\geqslant a.$ The latter contradicts with Lemma 2. So, $m_{k}$ may be chosen as an increasing sequence. Without loss of generality, going to renumbering, if required, we may consider that $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a$ holds for $m$ instead of $m_{k},$ i.e.,
$h(f_{m}(x_{m}),f_{m}(x^{\,\prime}_{m}))\geqslant a/2\qquad\forall\,\,m\in{\mathbb{N}}\,.$ (2.5)
We may assume that $x_{0}\neq\infty.$ Now, we use the condition 2): for every neighborhood $U$ of $x_{0}$ there is a neighborhood $V$ of $x_{0},$ $V\subset U,$ and $M=M(U,x_{0})\in{\mathbb{N}}$ such that $V\cap D_{m}$ is connected for every $m\geqslant M(U,x_{0}).$ Thus, given a ball $U_{1}=B(x_{0},2^{\,-1})$ there is $V_{1}\subset U_{1}$ and $M(1)\in{\mathbb{N}}$ such that $D_{m}\cap V_{1}$ is connected for $m\geqslant M(1).$ Since $x_{m},x_{m}^{\,\prime}\rightarrow x_{0}$ as $m\rightarrow\infty,$ there is $m_{1}\geqslant M(1)$ such that $x_{m_{1}},x^{\,\prime}_{m_{1}}\in V_{1}\cap D_{m_{1}}.$ Follow, given a ball $U_{2}=B(x_{0},2^{\,-2})$ there is $V_{2}\subset U_{2}$ and $M(2)\in{\mathbb{N}}$ such that $D_{m}\cap V_{2}$ is connected for $m\geqslant M(2).$ Since $x_{m},x_{m}^{\,\prime}\rightarrow x_{0}$ as $m\rightarrow\infty,$ there is $m_{2}\geqslant\max\{M(2),m_{1}\}$ such that $x_{m_{2}},x^{\,\prime}_{m_{2}}\in V_{2}\cap D_{m_{2}}.$ Etc. In general, given a ball $U_{k}=B(x_{0},2^{\,-k})$ there is $V_{k}\subset U_{k}$ and $M(k)\in{\mathbb{N}}$ such that $D_{m}\cap V_{k}$ is connected for $m\geqslant M(k).$ Since $x_{m},x_{m}^{\,\prime}\rightarrow x_{0}$ as $m\rightarrow\infty,$ there is $m_{k}\geqslant\max\{M(k),m_{k-1}\}$ such that $x_{m_{k}},x^{\,\prime}_{m_{k}}\in V_{k}\cap D_{m_{k}}.$ Relabeling, if required, we may consider that the same sequences $x_{m}$ and $x^{\prime}_{m}$ satisfy the above conditions, i.e., given a ball $U_{m}=B(x_{0},2^{\,-m}),$ $m=1,2,\ldots,$ there is $V_{m}\subset U_{m}$ such that $x_{m},x^{\,\prime}_{m}\in D_{m}\cap V_{m}$ while $D_{m}\cap V_{m}$ is connected. We join the points $x_{m}$ and $x^{\,\prime}_{m}$ by a path $\gamma_{m}:[0,1]\rightarrow D_{m}\cap V_{m}$ such that $\gamma_{m}(0)=x_{m},$ $\gamma_{m}(1)=x^{\,\prime}_{m}$ and $\gamma_{m}(t)\in D_{m}\cap V_{m}$ for $t\in(0,1),$ see Figure 1.
Figure 1: To prove Lemma 2
Denote by $C_{m}$ the image of the path $\gamma_{m}(t)$ under $f_{m}.$ It follows from relation (2.5) that
$h(C_{m})\geqslant a/2\qquad\forall\,m\in{\mathbb{N}}\,$, (2.6)
where $h$ denotes the chordal diameter of the set.
Let $\Gamma_{m}$ be a family of paths joining $\gamma_{m}$ and $A_{m}$ in $D_{m},$ where $A_{m}$ is a continuum from the condition 3). By the definition of the class $\mathfrak{H}^{p}_{Q,\delta}(\mathfrak{D}),$ the relations (1.3)–(1.4) hold at $x_{0}.$ Therefore,
$M_{p}(f_{m}(\Gamma_{m}))\leqslant\int\limits_{A(x_{0},\frac{1}{2^{m}},\varepsilon_{0})}Q(x)\cdot\eta^{p}(|x-x_{0}|)\ dm(x)$ (2.7)
for every Lebesgue measurable function $\eta:(\frac{1}{2^{m}},\varepsilon_{0})\rightarrow[0,\infty]$ such that $\int\limits_{\frac{1}{2^{m}}}^{\varepsilon_{0}}\eta(r)\,dr\geqslant 1.$ Observe that, the function
[…]
where $I(\varepsilon,\varepsilon_{0}):=\int\limits_{\varepsilon}^{\varepsilon_{0}}\psi(t)\,dt,$ satisfies the condition (1.4) with $r_{1}:=2^{\,-m},$ $r_{2}:=\varepsilon_{0}.$ Therefore, it follows from (2.4) and (2.7) that
$M_{p}(f_{m}(\Gamma_{m}))\leqslant\alpha(2^{\,-m})\rightarrow 0$ (2.8)
as $m\rightarrow\infty,$ where $\alpha(\varepsilon)$ is some nonnegative function tending to zero as $\varepsilon\rightarrow 0,$ which exists due to the condition (2.4).
On the other hand, observe that $f_{m}(\Gamma_{m})=\Gamma(C_{m},f_{m}(A_{m}),D_{m}^{\,\prime}).$ By the hypothesis of the lemma, $h(f_{m}(A_{m}))\geqslant\delta$ for all $m\in{\mathbb{N}}.$ Therefore, due to (2.6), $h(f_{m}(A_{m}))\geqslant\delta_{1}$ and $h(C_{m})\geqslant\delta_{1},$ where $\delta_{1}:=\min\{\delta,a/2\}.$ Using the fact that the domains $D_{m}^{\,\prime}:=f_{m}(D)$ are equi-uniform with respect to the $p$-modulus, we conclude that there exists $\sigma>0$ such that
$M_{p}(f_{m}(\Gamma_{m}))=M_{p}(\Gamma(C_{m},f_{m}(A_{m}),D_{m}^{\,\prime}))\geqslant\sigma\qquad\forall\,\,m\in{\mathbb{N}}\,$,
which contradicts the condition (2.8). The obtained contradiction indicates that the assumption of the absence of equicontinuity of the family $\mathfrak{H}^{p}_{Q,\delta}(\mathfrak{D})$ at $x_{0}$ was incorrect. The obtained contradiction completes the proof of the lemma. $\Box$
The following statement may be found in [Sev2, Lemmas 1.3, 1.4].
Proposition 2.1.   Let $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty],$ $n\geqslant 2,$ be a Lebesgue measurable function and let $x_{0}\in{\mathbb{R}}^{n}.$ Let $0<p\leqslant n.$ Assume that either of the following conditions holds
(a) $Q\in FMO(x_{0}),$
(b) $q_{x_{0}}(r)\,=\,O\left(\left[\log{\frac{1}{r}}\right]^{n-1}\right)$ as $r\rightarrow 0,$
(c) for some small $\delta_{0}=\delta_{0}(x_{0})>0$ we have the relations
$\int\limits_{\delta}^{\delta_{0}}\frac{dt}{t^{\frac{n-1}{p-1}}q_{x_{0}}^{\frac{1}{p-1}}(t)}<\infty,\qquad 0<\delta<\delta_{0},$ (2.9)
and
$\int\limits_{0}^{\delta_{0}}\frac{dt}{t^{\frac{n-1}{p-1}}q_{x_{0}}^{\frac{1}{p-1}}(t)}=\infty\,.$ (2.10)
Then there exist a number $\varepsilon_{0}\in(0,1)$ and a function $\psi(t)\geqslant 0$ satisfying (2.3) such that the relation (2.4) holds.
Uniting Lemma 2 and Proposition 2, we obtain the following.
Theorem 2.1.   Let $p\in(n-1,n],$ let $x_{0}\in\overline{\mathfrak{D}}$ and let $f_{m}\in\mathfrak{H}^{p}_{Q,\delta}(\mathfrak{D}),$ $m=1,2,\ldots,$ $f_{m}:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}$ be a sequence such that:
1) for every $m\in{\mathbb{N}},$ $D_{m}$ is locally connected at $\partial D_{m};$
2) for every neighborhood $U$ of $x_{0}$ there is a neighborhood $V$ of $x_{0},$ $V\subset U,$ and $M=M(U,x_{0})\in{\mathbb{N}}$ such that $V\cap D_{m}$ is connected for every $m\geqslant M=M(U,x_{0});$
3) there is $\varepsilon_{0}>0$ such that $A_{m}:=A_{f_{m}}\subset{\mathbb{R}}^{n}\setminus B(x_{0},\varepsilon_{0}),$ where $A_{f_{m}}$ is a continuum corresponding to the definition of the class $\mathfrak{H}^{p}_{Q,\delta}(\mathfrak{D})$ for $f_{m};$
4) the family $f_{m}(D_{m})$ is equi-uniform with respect to $p$-modulus over all $m\in{\mathbb{N}};$
5) at least one of two following conditions hold: $Q$ has a finite mean oscillation in $\overline{D_{0}},$ or (1.6) holds for every $x_{0}\in\overline{D_{0}}$ and some $\beta(x_{0})>0.$
Then the family $f_{m},$ $m=1,2,\ldots,$ is equicontinuous with respect to $x_{0},$ i.e., for any $\varepsilon>0$ there is $\delta=\delta(x_{0},\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(x^{\,\prime}))<\varepsilon$ whenever $x,x^{\,\prime}\in B(x_{0},\delta)\cap D_{m}$ and $m\in{\mathbb{N}}.$
3 On image of the continuum under homeomorphisms
The following statement holds, see, e.g., [Ku, Theorem 1.I.5.46]).
Proposition 3.1.   Let $A$ be a set in a topological space $X.$ If the set $C$ is connected and $C\cap A\neq\varnothing\neq C\setminus A,$ then $C\cap\partial A\neq\varnothing.$
Proposition 3.2.   Let $D$ and $G$ be domains in ${\mathbb{R}}^{n},$ $n\geqslant 2,$ such that $D\subset G,$ and let $K$ be a compactum in $D.$ Then $h(K,\partial D)\leqslant h(K,\partial G).$
Proof.   Since $\overline{{\mathbb{R}}^{n}}$ is connected, $\partial D\neq\varnothing.$ By the same reason, $\partial G\neq\varnothing.$ Thus, $h(K,\partial G)$ and $h(K,\partial D)$ are well-defined.
Let now $h(K,\partial D)=h(x_{0},y_{0}),$ $x_{0}\in K$ and $y_{0}\in\partial D,$ and let $h(K,\partial G)=h(z_{0},z_{0}),$ $z_{0}\in K$ and $w_{0}\in\partial G.$ We need to prove that $h(x_{0},y_{0})\leqslant h(z_{0},w_{0}).$ We prove this by the contradiction: assume that $h(z_{0},w_{0})<h(x_{0},y_{0}).$ Consider the ball $\overline{B_{h}(z_{0},h(z_{0},w_{0}))}.$ Since $D\subset G,$ the point $w_{0}$ cannot be inner point of $D.$ Let us join the points $z_{0}$ and $w_{0}$ by a path $\gamma$ in $\overline{B_{h}(z_{0},h(z_{0},w_{0}))}.$ Now, $|\gamma|\cap D\neq\varnothing\neq|\gamma|\setminus D.$ By Proposition 3 there is $v_{0}\in\partial D\cap|\gamma|.$ Thus, $v_{0}\in\overline{B_{h}(z_{0},h(z_{0},w_{0}))}$ that means $h(z_{0},v_{0})\leqslant h(z_{0},w_{0}).$ Since $z_{0}\in K$ and $v_{0}\in\partial D,$ the latter means that
$h(K,\partial D)\leqslant h(z_{0},v_{0})\leqslant h(z_{0},w_{0})<h(x_{0},y_{0})=h(K,\partial D)\,.$
The obtained contradiction completes the proof. $\Box$
Let $D$ be a domain in ${\mathbb{R}}^{n},$ $n\geqslant 2,$ and let $f:D\rightarrow{\mathbb{R}}^{n}$ (or $f:D\rightarrow\overline{{\mathbb{R}}^{n}})$ be a discrete open mapping, $\beta:[a,\,b)\rightarrow{\mathbb{R}}^{n}$ be a path, and $x\in\,f^{-1}(\beta(a)).$ A path $\alpha:[a,\,b)\rightarrow D$ is called a total $f$-lifting of $\beta$ starting at $x,$ if $(1)\quad\alpha(a)=x\,;$ $(2)\quad(f\circ\alpha)(t)=\beta(t)$ for any $t\in[a,b).$
Proposition 3.3.   Let $f:D\rightarrow{\mathbb{R}}^{n}$ be a discrete open and closed (boundary preserving) mapping, $\beta:[a,\,b)\rightarrow f(D)$ be a path, and $x\in\,f^{\,-1}\left(\beta(a)\right).$ Then $\beta$ has a total $f$-lifting starting at $x$ (see, e.g., [Vu, Lemma 3.7]).
The following statement holds, see e.g. [MRSY, Theorem 7.2].
Lemma 3.1. (Väisälä’s lemma on the weak flatness of inner points).   Let $n\geqslant 2$, let $D$ be a domain in $\overline{{\mathbb{R}}^{n}},$ and let $x_{0}\in D.$ Then for each $P>0$ and each neighborhood $U$ of point $x_{0}$ there is a neighborhood $V\subset U$ of the same point such that $M(\Gamma(E,F,D))>P$ for any continua $E,F\subset D$ intersecting $\partial U$ and $\partial V.$
[…]
Given numbers $\delta,r,N>0,$ a family $\mathfrak{D}$ of domains in ${\mathbb{R}}^{n},$ $n\geqslant 2,$ and a Lebesgue measurable function $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$, we denote by $\mathfrak{I}_{Q,\delta,N}(\mathfrak{D})$ a family of all open discrete and closed mappings $f:D\rightarrow\overline{{\mathbb{R}}^{n}}$ defined in some $D=D_{f}\in\mathfrak{D}$ and satisfying (1.3)–(1.4) at any point $x_{0}\in\overline{\mathfrak{D}}$ for all $0<r_{1}<r_{2}<\infty$ for which there exists a non-degenerate continuum $A=A_{f}$ and balls $\overline{B(x^{1}_{f},r^{1}_{f})},\ldots,\overline{B(x^{N}_{f},r^{N}_{f})},$ $x^{i}_{f}\in A_{f},$ $1\leqslant i\leqslant N,$ such that
$(1)\qquad A_{f}\subset\bigcup\limits_{i=1}^{N}\overline{B(x^{i}_{f},r^{i}_{f})}\subset D_{f}\,$,
$(2)\qquad h(\overline{B(x^{i}_{f},r^{i}_{f})},\partial D_{f})\geqslant\delta\,$,
$(3)\qquad h(f(A_{f}))\geqslant\delta$
and satisfying the condition: $h(E)\geqslant r$ whenever $E$ is any component of $\partial f(D_{f}).$ The following statement was proved for a fixed domain, see [RS2, Lemma 3]. Now we extend it for mappings defined on the family of domains.
Lemma 3.2.  Let $f_{m}\in\mathfrak{I}_{Q,\delta,N}(\mathfrak{D}).$ Assume that, for each point $x_{0}\in D_{0}$ and for every $0<r_{1}<r_{2}<r_{0}:=\sup\limits_{x\in D}|x-x_{0}|$ there is a set $E_{1}\subset[r_{1},r_{2}]$ of a positive linear Lebesgue measure such that the function $Q$ is integrable with respect to $\mathcal{H}^{n-1}$ over the spheres $S(x_{0},r)$ for every $r\in E_{1}.$ Then there exists $\delta_{1}>0$ such that $h(f_{m}(A_{m}),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots,$ where $A_{m}:=A_{f_{m}},$ $D_{m}:=D_{f_{m}}.$
The statement of Lemma 3 remains true if, instead of the condition regarding the integrability of the function $Q$ over spheres with respect to some set $E_{1}$ is replaced by a simpler condition: $Q\in L_{\rm loc}^{1}({\mathbb{R}}^{n}).$
Proof.   Since $D_{m}$ is a domain in ${\mathbb{R}}^{n},$ $m=1,2,\ldots,$ $\partial D_{m}\neq\varnothing.$ In addition, since $C(\partial D_{m},f_{m})\subset\partial f_{m}(D_{m})$ whenever $f_{m}$ is open, discrete and closed (see [Vu, Theorem 3.3]), $\partial f_{m}(D_{m})\neq\varnothing,$ as well. Thus, the quantity $h(f_{m}(A_{m}),\partial f_{m}(D_{m}))$ is well-defined, so the formulation of the lemma is correct.
Let us prove Lemma 3 by the contradiction, partially using the approach of the proof of Lemma 4.1 in [SevSkv]. Suppose that the conclusion of the lemma is not true. Then for each $k\in{\mathbb{N}}$ there is some number $m=m_{k}$ such that $h(f_{m_{k}}(A_{m_{k}}),\partial f_{m_{k}}(D_{m_{k}}))<1/k,$ $A_{m}:=A_{f_{m}}.$ In order not to complicate the notation, we will further assume that $h(f_{m}(A_{m}),\partial f_{m}(D_{m}))<1/m,$ $m=1,2,\ldots.$ Note that the set $f_{m}(A_{m})$ is compact as a continuous image of a compact set $A_{m}\subset D_{m}$ under the mapping $f_{m}.$ In this case, there are elements $x_{m}\in f_{m}(A_{m})$ and $y_{m}\in\partial f_{m}(D_{m})$ such that $h(f_{m}(A_{m}),\partial f_{m}(D_{m}))=h(x_{m},y_{m})<1/m.$ Let $E_{m}$ be a component of $\partial f_{m}(D_{m})$ containing $y_{m}.$ Due to the compactness of $\overline{{\mathbb{R}}^{n}},$ we may assume that $y_{m}\rightarrow y_{0}$ as $m\rightarrow\infty;$ then also $x_{m}\rightarrow y_{0}$ as $m\rightarrow\infty.$
Put $P>0$ and $U=B_{h}(y_{0},r_{0})=\{y\in\overline{{\mathbb{R}}^{n}}:h(y,y_{0})<r_{0}\},$ where $2r_{0}:=\min\{r/2,\delta/2\},$ where $r$ and $\delta$ are numbers from the condition of the lemma. Observe that $E_{m}\cap U\neq\varnothing\neq E_{m}\setminus U$ for sufficiently large $m\in{\mathbb{N}},$ since $y_{m}\rightarrow y_{0}$ as $m\rightarrow\infty,$ $y_{m}\in E_{m};$ besides that $h(E_{m})\geqslant r>r/2\geqslant 2r_{0}$ and $h(U)\leqslant 2r_{0}.$ Observe that, $E_{m}$ is closed (see [Ku, Theorem 1.III.5.46]). Now, $E_{m}$ is a continuum, so that $E_{m}\cap\partial U\neq\varnothing$ by Proposition 3. Similarly, $f_{m}(A_{m})\cap U\neq\varnothing\neq f_{m}(A_{m})\setminus U$ for sufficiently large $m\in{\mathbb{N}},$ since $x_{m}\rightarrow y_{0}$ as $m\rightarrow\infty,$ $x_{m}\in f_{m}(A_{m});$ besides that $h(f_{m}(A_{m}))\geqslant\delta>\delta/2\geqslant 2r_{0}$ and $h(U)\leqslant 2r_{0}.$ Since $f_{m}(A_{m})$ is a continuum, $f_{m}(A_{m})\cap\partial U\neq\varnothing$ by Proposition 3. By the proving above,
$E_{m}\cap\partial U\neq\varnothing\neq f_{m}(A_{m})\cap\partial U$ (3.1)
for sufficiently large $m\in{\mathbb{N}}.$ By Lemma 3 there is $\subset U, VV$ is a neighborhood of $y_{0},$ such that
$M(\Gamma(E,F,\overline{{\mathbb{R}}^{n}}))>P$ (3.2)
for any continua $E,F\subset\overline{{\mathbb{R}}^{n}}$ with $E\cap\partial U\neq\varnothing\neq E\cap\partial V$ and $F\cap\partial U\neq\varnothing\neq F\cap\partial V.$ Arguing similarly to above, we may prove that
$E_{m}\cap\partial V\neq\varnothing\neq f_{m}(A_{m})\cap\partial V$
for sufficiently large $m\in{\mathbb{N}}.$ Thus, by (3.2)
$M(\Gamma(f_{m}(A_{m}),E_{m},\overline{{\mathbb{R}}^{n}}))>P$ (3.3)
for sufficiently large $m=1,2,\ldots.$ We now prove that the relation (3.3) contradicts the definition of $f_{m}$ in (1.3)–(1.4).
Indeed, let $\gamma:[0,1]\rightarrow\overline{{\mathbb{R}}^{n}}$ be a path in $\Gamma(f_{m}(A_{m}),E_{m},\overline{{\mathbb{R}}^{n}}),$ i.e., $\gamma(0)\in f_{m}(A_{m}),$ $\gamma(1)\in E_{m}$ and $\gamma(t)\in\overline{{\mathbb{R}}^{n}}$ for $t\in(0,1).$ Let $t_{m}=\sup\limits_{\gamma(t)\in f_{m}(D_{m})}t$ and let $\alpha_{m}(t)=\gamma|_{[0,t_{m})}.$ Let $\Gamma_{m}$ consists of all such paths $\alpha_{m},$ now $\Gamma(f_{m}(A_{m}),E_{m},\overline{{\mathbb{R}}^{n}})>\Gamma_{m}$ and by the minorization principle of the modulus (see [Fu, Theorem 1])
$M(\Gamma_{m})\geqslant M(\Gamma(f_{m}(A_{m}),E_{m},\overline{{\mathbb{R}}^{n}}))\,.$ (3.4)
Let $\Delta_{m}$ be a family of all total $f_{m}$ liftings of $\Gamma_{m}$ starting at $A_{m}$ (it exists by Proposition 3). Now
$M(f_{m}(\Delta_{m}))=M(\Gamma_{m})\,.$ (3.5)
Let $\beta_{m}:[0,t_{m})\rightarrow D_{m},$ $\beta_{m}\in\Delta_{m},$ $f_{m}\circ\beta_{m}=\alpha_{m}.$ Since open discrete and closed mappings preserve the boundary ([Vu, Theorem 3.3]), $\beta_{m}(t)\rightarrow\partial D_{m}$ as $t\rightarrow t_{m}.$
Let $A_{m}\subset\bigcup\limits_{i=1}^{N}\overline{B(z_{mi},r_{mi})},$ $r_{mi}>0,$ where $z_{mi}\in A_{m},$ $1\leqslant i\leqslant N$ and $h(\overline{B(z_{mi},r_{mi})},\partial D_{m})\geqslant\delta.$ By the compactness of $\overline{{\mathbb{R}}^{n}},$ we may consider that $z_{mi}\rightarrow z_{0i}$ as $m\rightarrow\infty.$
Since $h(\overline{B(z_{mi},r_{mi})},\partial D_{m})\geqslant\delta,$ we have that $B_{h}(z_{mi},\delta/2)\subset D_{m}$ for every $m=1,2,\ldots$ and $1\leqslant i\leqslant N,$ where $B_{h}(w_{0},p):=\{x\in\overline{{\mathbb{R}}^{n}}:h(x,w_{0})<p\}.$ Indeed, otherwise $B_{h}(z_{mi},\delta/2)\cap D_{m}\neq\varnothing\neq B_{h}(z_{mi},\delta/2)\setminus D_{m}.$ Since $B_{h}(z_{mi},\delta/2)$ is connected, we have that $B_{h}(z_{mi},\delta/2)\cap\partial D_{m}\neq\varnothing.$ Now, $h(z_{mi},\partial D_{m})\leqslant h(\overline{B(z_{mi},r_{mi})},\partial D_{m})\leqslant\delta/2,$ which contradicts with the condition $h(\overline{B(z_{mi},r_{mi})},\partial D_{m})\geqslant\delta.$ Thus, $h(\overline{B(z_{mi},r_{mi})},\partial D_{m})\geqslant\delta,$ as required.
Since $z_{mi}\rightarrow z_{0i}$ as $m\rightarrow\infty,$ we have that $h(z_{mi},z_{0i})<\delta/4$ for sufficiently large $m\in{\mathbb{N}}.$ Now, $z_{0i}\in B_{h}(z_{mi},\delta/2)\subset D_{m}$ for above $m.$ In particular, it follows from the mentioned above that $z_{0i}\neq\infty.$ We may consider that, $z_{0i}\in D_{m}$ for all $m\in{\mathbb{N}}.$
In the similar way, we may consider that $r_{mi}\rightarrow r_{0i}$ as $m\rightarrow\infty$ for some $r_{0i}\in{\mathbb{R}}^{\,+}\cup\{\infty\}.$ Observe that $r_{0i}\neq\infty.$ Otherwise, there is $w_{m}\in B(z_{mi},r_{mi}),$ $w_{m}\in S(z_{mi},r_{mi}/2),$ where $z_{mi}\rightarrow z_{0i}\neq\infty$ as $m\rightarrow\infty,$ and $r_{mi}\rightarrow\infty$ as $m\rightarrow\infty.$ Now, by the triangle inequality,
$|w_{m}|\geqslant|w_{m}-z_{mi}|-|z_{mi}|\geqslant r_{mi}/2-c\rightarrow\infty\,$,
$m\rightarrow\infty,$ where $c>0$ is some constant. Thus, $w_{m}\in B(z_{mi},r_{mi})$ and $w_{m}\rightarrow\infty$ as $m\rightarrow\infty.$ Now, $h(\overline{B(z_{mi},r_{mi})},\infty)\rightarrow 0$ as $m\rightarrow\infty.$ The latter contradicts the condition $h(\overline{B(z_{mi},r_{mi})},\partial D_{m})\geqslant\delta.$ Indeed, by Proposition 3
$h(\overline{B(z_{mi},r_{mi})},\infty)\geqslant h(\overline{B(z_{mi},r_{mi})},\partial D_{m})\geqslant\delta\,.$
The latter proves that $r_{0i}\neq\infty.$
Let $x\in\overline{B(z_{mi},r_{mi})},$ $r_{mi}\geqslant 0.$ Now, by the triangle inequality
$|x-z_{0i}|\leqslant|x-z_{mi}|+|z_{mi}-z_{0i}|<r_{0i}+\delta_{mi}\,$, (3.6)
where $\delta_{mi}>0,$ $\delta_{mi}\rightarrow 0$ as $m\rightarrow\infty.$ The relation (3.6) implies that
$\overline{B(z_{mi},r_{mi})}\subset B(z_{0i},r_{0i}+\delta_{mi})\,.$ (3.7)
Observe that,
$\overline{B(z_{0i},r_{0i})}\subset D_{m}$ (3.8)
for sufficiently large $m\in{\mathbb{N}}.$ Indeed, let $\alpha_{mi}=h(\overline{B(z_{0i},r_{0i})},\partial D_{m})=h(a_{mi},b_{mi}),$ where $a_{mi}\in\overline{B(z_{0i},r_{0i})}$ and $b_{mi}\in\partial D_{m}.$ Let $h(\overline{B(z_{mi},r_{mi})},\partial D_{m})=h(c_{mi},d_{mi})\geqslant\delta>0,$ $c_{mi}\in\overline{B(z_{mi},r_{mi})},$ $d_{mi}\in\partial D_{m}.$ Now, by the triangle inequality
$\alpha_{mi}=h(a_{mi},b_{mi})\geqslant h(b_{mi},c_{mi})-h(c_{mi},a_{mi})\geqslant$
$\geqslant\delta-|c_{mi}-a_{mi}|\geqslant\delta-|c_{mi}-z_{mi}|-|z_{mi}-z_{0i}|-|z_{0i}-a_{mi}|$ (3.9)
$\geqslant\delta-r_{mi}-r_{0i}-|z_{mi}-z_{0i}|\geqslant\delta/2$
for sufficiently large $m\in{\mathbb{N}}.$ Now, (3.9) proves (3.8). Since $\delta_{mi}\rightarrow 0$ as $m\rightarrow\infty$ in (3.6), the relation (3.8) implies that we may find some $r^{i}_{2}>r^{i}_{1}>r_{0i}>0$ such that $r_{0i}+\delta_{mi}<r^{i}_{1}$ and simultaneously $\overline{B(z_{0i},r^{i}_{2})}\subset D_{m}$ for sufficiently large $m.$ Now, (3.7) implies that $A_{m}\subset\bigcup\limits_{i=1}^{N}B(z_{0i},r^{i}_{1}).$ We now show that
$\Delta_{m}>\bigcup\limits_{i=1}^{N}\Gamma(S(z_{0i},r^{i}_{1}),S(z_{0i},r^{i}_{2}),A(z_{0i},r^{i}_{1},r^{i}_{2}))\,.$ (3.10)
Indeed, let $\gamma\in\Delta_{m},$ in other words, $\gamma:[0,t_{m})\rightarrow D_{m},$ $\gamma(0)\in A_{m}$ and $\gamma(t)\rightarrow\partial D_{m}$ as $t\rightarrow t_{m}-0.$ Now, there is $1\leqslant i\leqslant N$ such that $|\gamma|\cap B(z_{0i},r^{i}_{1})\neq\varnothing\neq|\gamma|\cap(D_{m}\setminus B(z_{0i},r^{i}_{1})).$ Therefore, by Proposition 3 there is $0<t^{\,*}_{1}<t_{m}$ such that $\gamma(t^{\,*}_{1})\in S(z_{0i},r^{i}_{1}).$ We may assume that $\gamma(t)\not\in B(z_{0i},r^{i}_{1})$ for $t>t^{\,*}_{1}.$ Put $\gamma_{1}:=\gamma|_{[t^{\,*}_{1},t_{m}]}.$ Similarly, $|\gamma_{1}|\cap B(z_{0i},r^{i}_{2})\neq\varnothing\neq|\gamma_{1}|\cap(D_{m}\setminus B(z_{0i},r^{i}_{2})).$ By Proposition 3 there is $t^{\,*}_{1}<t^{\,*}_{2}<t_{m}$ with $\gamma(t^{\,*}_{2})\in S(z_{0i},r^{i}_{2}).$ We may assume that $\gamma(t)\in B(z_{0i},r^{i}_{2})$ for $t<t^{\,*}_{2}.$ Put $\gamma_{2}:=\gamma|_{[t^{\,*}_{1},t^{\,*}_{2}]}.$ Then, the path $\gamma_{2}$ is a subpath of $\gamma,$ which belongs to the family $\Gamma(S(z_{0i},r^{i}_{1}),S(z_{0i},r^{i}_{2}),A(z_{0i},r^{i}_{1},r^{i}_{2})).$ Thus, the relation (3.10) is established.
On the other hand, we set
$\psi_{i}(t)=\left\{\begin{array}[]{rr}1/[tq_{z_{0i}}^{\frac{1}{n-1}}(t)]\ ,&\ t\in(r^{i}_{1},r^{i}_{2})\ ,\\ 0\ ,&\ t\notin(r^{i}_{1},r^{i}_{2})\ .\end{array}\right$.
Now,
$\int\limits_{A(z_{0i},r^{i}_{1},r^{i}_{2})}Q(x)\cdot\psi^{n}(|x-z_{0i}|)\,dm(x)=\omega_{n-1}I_{i}\,$, (3.11)
where $I_{i}=\int\limits_{r^{i}_{1}}^{r^{i}_{2}}\frac{dt}{tq_{z_{0i}}^{\frac{1}{n-1}}(t)}.$ Note that, the function $\eta_{1}(t)=\psi(t)/I_{i},$ $t\in(r^{i}_{1},r^{i}_{2}),$ satisfies (1.4), because $\int\limits_{r^{i}_{1}}^{r^{i}_{2}}\eta_{1}(t)\,dt=1.$ Now, it follows from (3.11) that
$M(f_{m}(\Gamma(S(z_{0i},r^{i}_{1}),S(z_{0i},r^{i}_{2})),A(z_{0i},r^{i}_{1},r^{i}_{2})))\leqslant\frac{\omega_{n-1}}{I_{i}^{n-1}}<\infty\,$, (3.12)
where $I_{i}=I_{i}(z_{0i},r^{i}_{1},r^{i}_{2})=\int\limits_{r^{i}_{1}}^{r^{i}_{2}}\frac{dr}{rq_{z_{0i}}^{\frac{1}{n-1}}(r)}$ and $I_{i}<\infty$ because $q_{z_{0i}}(r)<\infty$ for $r\in E_{1}$ and some set of positive linear Lebesgue measure by the assumption of the lemma.
Finally, by (3.4), (3.5), (3.10) and (3.12) we obtain that
$M(\Gamma(f_{m}(A_{m}),E_{m},\overline{{\mathbb{R}}^{n}}))\leqslant M(\Gamma_{m})$
$=M(f_{m}(\Delta_{m}))\leqslant\sum\limits_{i=1}^{n}M(f_{m}(\Gamma(S(z_{0i},r^{i}_{1}),S(z_{0i},r^{i}_{2}),A(z_{0i},r^{i}_{1},r^{i}_{2}))))$ (3.13)
$\leqslant\sum\limits_{i=1}^{n}\frac{\omega_{n-1}}{I_{i}^{n-1}}:=C<\infty\,.$
Since $P$ in (3.3) may be done arbitrary big, the relations (3.3) and (3.13) contradict each other. This completes the proof. $\Box$
4 Proof of the main results
Proof of Theorem 1. Let us prove the theorem by contradiction. Assume the assertion of Theorem 1 is false. Then for some $\varepsilon_{0}>0$ and for numbers $\delta_{k}=\frac{1}{k},$ $k=1,2,\ldots,$ there are $m_{k}\in{\mathbb{N}}$ and $x_{k},y_{k}\in D_{m_{k}}$ such that $h(f_{m_{k}}(x_{k}),f_{m_{k}}(y_{k}))\geqslant\varepsilon_{0}$ whenever $|x_{k}-y_{k}|<\frac{1}{k}.$ By Lemma 2 every $f_{m_{k}}$ has a continuous extension to $\partial D_{m_{k}},$ so that $f_{m_{k}}$ is uniformly continuous in $\overline{D_{m_{k}}}.$ Thus, we may consider that the sequence $m_{k},$ $k=1,2,\ldots,$ is increasing. Without loss of generality, we may consider that the above holds for $m=1,2,\ldots,$ but not for $m_{k}.$ In particular,
$h(f_{m}(x_{m}),f_{m}(y_{m}))\geqslant\varepsilon_{0}\,,\qquad m\in{\mathbb{N}}\,.$ (4.1)
Now, by the triangle inequality,
$h(f_{m}(x_{m}),f_{m}(y_{m}))\leqslant h(f_{m}(x_{m}),f_{m}(x_{0}))+h(f_{m}(x_{0}),f_{m}(y_{m}))\,.$ (4.2)
By the compactness of $\overline{{\mathbb{R}}^{n}},$ we may consider that $x_{m},y_{m}\rightarrow x_{0}$ as $m\rightarrow\infty.$ There are two cases: 1) $x_{0}\in D_{0},$ 2) $x_{0}\in\partial D_{0}.$ In the first case, when $x_{0}$ is the inner point of $D_{0},$ all mappings $f_{m}$ except of finite numbers are defined in some ball $B(x_{0},\varepsilon_{0})\subset D_{0}\cap D_{m},$ $m\geqslant M_{0}.$ Now, $f_{m}$ is equicontinuous at $x_{0}$ (see [RS1, Theorems 6.1, 6.5] for $p=n$ and [Sev2, Theorems 4.1, 4.3] for n1<p<n.n-1<p<n. In this case, the right side in (4.2) tends to zero as $m\rightarrow\infty$ which contradicts with (4.1), and the proof finishes.
Let us consider the case 2) $x_{0}\in\partial D_{0}.$ Observe that $A_{m}:=A\subset{\mathbb{R}}^{n}\setminus B(x_{0},\varepsilon_{0})$ for some $\varepsilon_{0}>0,$ because the $A$ is a continuum in $D_{0}.$ Thus, all of conditions of Theorem 2 are satisfied. In particular, it follows from this theorem that $f_{m}$ has a continuous extension to $\partial D_{m}$ for any $m\in{\mathbb{N}}.$ Thus, there is $x_{m}\in D_{m}$ such that $h(f_{m}(x^{\,\prime}_{m}),f_{m}(x_{0}))<\frac{1}{m}.$ Now, by the triangle inequality, by (4.2) and by Theorem 2 we obtain that
$h(f_{m}(x_{m}),f_{m}(y_{m}))\leqslant h(f_{m}(x_{m}),f_{m}(x^{\,\prime}_{m}))+h(f_{m}(x^{\,\prime}_{m}),f_{m}(x_{0}))+$
$+h(f_{m}(x_{0}),f_{m}(x^{\,\prime}_{m}))+h(f_{m}(x^{\,\prime}_{m}),f_{m}(y_{m}))<$ (4.3)
$+h(f_{m}(x_{m}),f_{m}(x^{\,\prime}_{m}))+h(f_{m}(x^{\,\prime}_{m}),f_{m}(y_{m}))+\frac{2}{m}\rightarrow 0\qquad\text{as}\qquad m\rightarrow\infty.$
The relations (4.1) and (4.3) contradict each other and prove the uniformly equicontinuity of $f_{m}.$
Given a compactum $K$ in $D_{0},$ there is $m_{0}=m_{0}(K)$ such that $K\subset D_{m}$ for $m\geqslant m_{0}(K).$ Now, by the proving above, $f_{m},$ $m=1,2,\ldots,$ is uniformly equicontinuous family on $K.$ By Arzela-Ascoli Theorem (see. e.g., [Va, Theorem 20.4]) $f_{m}$ forms a normal family on $K,$ i.e., there exists a subsequence $f_{m_{k}},$ $k=1,2,\ldots,$ such that $f_{m_{k}}(x)\rightarrow f(x)$ as $k\rightarrow\infty.$ Since $K$ is an arbitrary compactum in $D_{0},$ the mapping $f(x)$ is defined in all $D_{0}.$ Now, we prove that $f$ has a continuous extension at every point $x_{0}\in\partial D_{0}.$ By the proving above, given $\varepsilon>0$ there is $\delta=\delta(\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(y))<\varepsilon$ whenever $x,y\in D_{m},$ $|x-y|<\delta$ and $m\in{\mathbb{N}}.$ Fix $x_{0}\in\partial D_{0}.$ Let $\delta_{*}=\delta_{*}(\varepsilon)=\delta/2$ and let $x,y\in B(x_{0},\delta_{*})\cap D_{0}.$ Now, by the triangle inequality $|x-y|<2\delta_{*}=\delta.$ Since $x,y\in D_{0},$ there is $m_{0}=m_{0}(x,y)\in{\mathbb{N}}$ such that $x,y\in D_{0}$ for every $m\geqslant m_{0}.$ Taking into account the mentioned above, we obtain that $h(f_{m}(x),f_{m}(y))<\varepsilon,$ $m\geqslant m_{0}.$ Taking the limit as $m\rightarrow\infty$ in the latter inequality, we obtain that $h(f(x),f(y))\leqslant\varepsilon.$ Now, $f$ satisfies the Cauchy condition at $x_{0},$ so that $f$ has a limit at $x_{0},$ as required.
It remains to show that $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots$ and some $\delta_{1}>0$ under conditions $p=n$ and 5), 6) of the theorem. We cover the continuum $A$ by the balls $B(x,r)$ with $x\in A$ and $0<r<\frac{1}{2}{\rm dist}\,(A,\partial D_{0}).$ Since $A$ is a compactum, we may choose a finite subcover $\bigcup\limits_{i=1}^{N}B(x_{i},r),$ $x_{i}\in A,$ of $A.$ Observe that, $h(\overline{B(x_{i},r)},\partial D_{m})\geqslant r_{*}>0$ for some $r_{*}>0,$ all $1\leqslant i\leqslant N$ and all $m\in{\mathbb{N}}$ except finite numbers. Otherwise, there are sequences $x_{m}\in\overline{B(x_{i},r)}$ and $y_{m}\in\partial D_{m}$ with $h(x_{m},y_{m})\rightarrow 0$ as $m\rightarrow\infty.$ We may consider that $x_{m},y_{m}\rightarrow x_{0}$ as $m\rightarrow\infty.$ Obviously, $x_{0}\in\overline{B(x_{i},r)}$ by the closeness of $\overline{B(x_{i},r)}.$ Now, there is $\varepsilon_{0}>0$ such that $\overline{B(x_{0},\varepsilon_{0})}\subset D_{0}$ and $\overline{B(x_{0},\varepsilon_{0})}$ belongs to all $D_{m}$ except of finite numbers. However, the latter contradicts the condition $y_{m}\rightarrow x_{0}$ as $m\rightarrow\infty$ while $y_{m}\in\partial D_{m}.$ Indeed, in this case $\overline{B(x_{0},\varepsilon_{0})}\cap\partial D_{m}\neq\varnothing$ for infinitely many $m\in{\mathbb{N}}.$
Finally, observe that $f_{m}\in\mathfrak{I}_{Q,\delta,N}(\mathfrak{D}),$ where $\mathfrak{D}=\{D_{m}\}_{m=1}^{\infty},$ $A_{f_{m}}=A.$ By Lemma 3, $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots$ and some $\delta_{1}>0,$ as required. $\Box$
5 Main Lemma for mappings with a branching
Given $p\geqslant 1,$ a number $\delta>0,$ a family $\mathfrak{D}$ of domains in ${\mathbb{R}}^{n},$ $n\geqslant 2,$ and a Lebesgue measurable function $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$, we denote by $\mathfrak{J}_{Q,\delta,p,E}(\mathfrak{D})$ the family of all open discrete closed ring mappings $f:D_{f}\rightarrow\overline{{\mathbb{R}}^{n}}\setminus E$ defined in some a domain $D_{f}\in\mathfrak{D}$ and satisfying (1.3)–(1.4) at any point $x_{0}\in\overline{\mathfrak{D}}$ such that the following condition holds: for any domain $D^{\,\prime}_{f}:=f(D_{f})$ there exists a continuum $K_{f}\subset D^{\,\prime}_{f}$ such that $h(K_{f})\geqslant\delta$ and $h(f^{\,-1}(K_{f}),\partial D_{f})\geqslant\delta>0.$ The following statement holds.
The following lemma is an analog of Lemma 2 for mappings which are not homeomorphisms.
Lemma 5.1.   Let $p\in(n-1,n],$ let $x_{0}\in\overline{\mathfrak{D}}$ and let $f_{m}\in\mathfrak{J}_{Q,\delta,p,E}(\mathfrak{D}),$ $m=1,2,\ldots,$ $f_{m}:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}$ be a sequence such that:
1) for every $m\in{\mathbb{N}},$ a domain $D_{m}$ is locally connected at $\partial D_{m};$
2) for every neighborhood $U$ of $x_{0}$ there is a neighborhood $V$ of $x_{0},$ $V\subset U,$ and $M=M(U,x_{0})\in{\mathbb{N}}$ such that $V\cap D_{m}$ is connected for every $m\geqslant M(U,x_{0});$
3) there is $\varepsilon_{0}>0$ such that $f^{\,-1}_{m}(K_{m})\subset{\mathbb{R}}^{n}\setminus B(x_{0},\varepsilon_{0}),$ where $K_{m}:=K_{f_{m}}$ is a continuum corresponding to the definition of the class $\mathfrak{J}_{Q,\delta,p,E}(\mathfrak{D})$ for $f_{m};$
4) the family $f_{m}(D_{m})$ is equi-uniform with respect to $p$-modulus over all $m\in{\mathbb{N}};$
5) there exists $\varepsilon_{0}=\varepsilon_{0}(x_{0})>0$ and a Lebesgue measurable function $\psi:(0,\varepsilon_{0})\rightarrow[0,\infty]$ such that relations (2.3)–(2.4) hold as $\varepsilon\rightarrow 0,$ where $A(x_{0},\varepsilon,\varepsilon_{0})$ is defined in (1.2).
Then the family $f_{m},$ $m=1,2,\ldots,$ is equicontinuous with respect to $x_{0},$ i.e., for any $\varepsilon>0$ there is $\delta=\delta(x_{0},\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(x^{\,\prime}))<\varepsilon$ whenever $x,x^{\,\prime}\in B(x_{0},\delta)\cap D_{m}$ and $m\in{\mathbb{N}}.$
Proof.   Suppose the contrary. Then there exists $x_{0}\in\overline{\mathfrak{D}}$ and a number $a>0$ with the following property: for $\delta=1/k,$ $k=1,2,\ldots,$ there are $m_{k}\in{\mathbb{N}}$ and $x_{k},x^{\,\prime}_{k}\in B(x_{0},1/k)\cap D_{m_{k}}$ such that $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a.$ Arguing similarly to the proof of Lemma 2, we may consider that the sequence $m_{k}$ is increasing by $k.$ Without loss of generality, going to renumbering, if required, we may consider that $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a$ holds for $m$ instead of $m_{k},$ i.e., (2.5) is satisfied. We may assume that $x_{0}\neq\infty.$ Now, we use the condition 2): for every neighborhood $U$ of $x_{0}$ there is a neighborhood $V$ of $x_{0},$ $V\subset U,$ and $M=M(k)$ such that $V\cap D_{m}$ is connected for every $m\geqslant M(k).$ Thus, given a ball $U_{1}=B(x_{0},2^{\,-1})$ there is $V_{1}\subset U_{1}$ and $M(1)\in{\mathbb{N}}$ such that $D_{m}\cap V_{1}$ is connected for $m\geqslant M(1).$ Since $x_{m},x_{m}^{\,\prime}\rightarrow x_{0}$ as $m\rightarrow\infty,$ there is $m_{1}\geqslant M(1)$ such that $x_{m_{1}},x^{\,\prime}_{m_{1}}\in V_{1}\cap D_{m_{1}}.$ Follow, given a ball $U_{2}=B(x_{0},2^{\,-2})$ there is $V_{2}\subset U_{2}$ and $M(2)\in{\mathbb{N}}$ such that $D_{m}\cap V_{2}$ is connected for $m\geqslant M(2).$ Since $x_{m},x_{m}^{\,\prime}\rightarrow x_{0}$ as $m\rightarrow\infty,$ there is $m_{2}\geqslant\max\{M(2),m_{1}\}$ such that $x_{m_{2}},x^{\,\prime}_{m_{2}}\in V_{2}\cap D_{m_{2}}.$ Etc. In general, given a ball $U_{k}=B(x_{0},2^{\,-k})$ there is $V_{k}\subset U_{k}$ and $M(k)\in{\mathbb{N}}$ such that $D_{m}\cap V_{k}$ is connected for $m\geqslant M(k).$ Since $x_{m},x_{m}^{\,\prime}\rightarrow x_{0}$ as $m\rightarrow\infty,$ there is $m_{k}\geqslant\max\{M(k),m_{k-1}\}$ such that $x_{m_{k}},x^{\,\prime}_{m_{k}}\in V_{k}\cap D_{m_{k}}.$ Relabeling, if required, we may consider that the same sequences $x_{m}$ and $x^{\prime}_{m}$ satisfy the above conditions, i.e., given a ball $U_{m}=B(x_{0},2^{\,-m}),$ $m=1,2,\ldots,$ there is $V_{m}\subset U_{m}$ such that $x_{m},x^{\,\prime}_{m}\in D_{m}\cap V_{m}$ while $D_{m}\cap V_{m}$ is connected. We join the points $x_{m}$ and $x^{\,\prime}_{m}$ by a path $\gamma_{m}:[0,1]\rightarrow D_{m}\cap V_{m}$ such that $\gamma_{m}(0)=x_{m},$ $\gamma_{m}(1)=x^{\,\prime}_{m}$ and $\gamma_{m}(t)\in D_{m}\cap V_{m}$ for $t\in(0,1),$ see Figure 2.
Figure 2: To prove Lemma 5
Denote by $C_{m}$ the image of the path $\gamma_{m}$ under the mapping $f_{m}.$ It follows from relation (2.5) that a condition of the form (2.6) is satisfied, where $h$ denotes the chordal diameter of the set.
By the definition of the family $\mathfrak{J}_{Q,\delta,p,E}(\mathfrak{D}),$ for any $f_{m}\in\mathfrak{J}_{Q,\delta,p,E}(\mathfrak{D})$ and any domain $D^{\,\prime}_{m}:=f_{m}(D_{m})$ there exists a continuum $K_{m}\subset D^{\,\prime}_{m}$ such that $h(K_{m})\geqslant\delta$ and $h(f_{m}^{\,-1}(K_{m}),\partial D_{m})\geqslant\delta>0.$ Since by the condition of the lemma the domains $D^{\,\prime}_{m}$ form the equi-uniform family with respect to $p$-modulus and due to (2.6), we obtain that the inequality
$M_{p}(\Gamma(K_{m},C_{m},D^{\,\prime}_{m}))\geqslant b$ (5.1)
holds for all $m=1,2,\ldots$ and some $b>0$ Consider the family $\Gamma_{m},$ consisting of all paths $\beta:[0,1)\rightarrow D^{\,\prime}_{m},$ where $\beta(0)\in C_{m}$ and $\beta(t)\rightarrow p\in K_{m}$ for $t\rightarrow 1.$ Let $\Gamma^{*}_{m}$ be a family of all total $f_{m}$-liftings $\alpha:[0,1)\rightarrow D$ of the family $\Gamma_{m}$ with origin on $\gamma_{m}$ (see Proposition 3). Since $f_{m}$ is closed, we have: $\alpha(t)\rightarrow f^{\,-1}_{m}(K_{m}),$ where $f^{\,-1}_{m}(K_{m})$ is the total preimage of the continuum $K_{m}$ under the mapping $f_{m}.$ By (1.3), it follows that
$M_{p}(f_{m}(\Gamma_{m}^{*}))$
$\leqslant M_{p}(f_{m}(\Gamma(|\gamma_{m}|,f^{\,-1}_{m}(K_{m}),D_{m})))\leqslant\int\limits_{A(x_{0},\frac{1}{2^{m}},\varepsilon_{0})}Q(x)\cdot\eta^{p}(|x-x_{0}|)\ dm(x)$ (5.2)
for each measurable function $\eta:(\frac{1}{2^{m}},\varepsilon_{0})\rightarrow[0,\infty],$ such that $\int\limits_{\frac{1}{2^{m}}}^{\varepsilon_{0}}\eta(r)\,dr\geqslant 1.$ Note that, the function
[…]
where $I(\varepsilon,\varepsilon_{0}):=\int\limits_{\varepsilon}^{\varepsilon_{0}}\psi(t)\,dt,$ satisfies the normalization condition (1.4) with $r_{1}:=2^{\,-m},$ $r_{2}:=\varepsilon_{0}.$ Therefore from the conditions (2.3)–(2.4) and (5.2) it follows that
$M_{p}(f_{m}(\Gamma^{\,*}_{m}))\leqslant\alpha(2^{\,-m})\rightarrow 0$ (5.3)
as $m\rightarrow\infty,$ where $\alpha(\varepsilon)$ is some nonnegative function tending to zero as $\varepsilon\rightarrow 0,$ which exists by the condition (2.4). In addition, $f_{m}(\Gamma^{\,*}_{m})=\Gamma_{m}$ and $M_{p}(\Gamma_{m})=M_{p}(\Gamma(K_{m},C_{m},D^{\,\prime}_{m})).$ Thus
$M_{p}(f_{m}(\Gamma^{\,*}_{m}))=M_{p}(\Gamma(K_{m},C_{m},D^{\,\prime}_{m}))\,.$ (5.4)
However, relations (5.3) and (5.4) contradict with (5.1). The resulting contradiction indicates that the assumption (2.5) was incorrect, and, therefore, the family $\mathfrak{J}_{Q,\delta,p,E}(\mathfrak{D})$ is equicontinuous at $x_{0}\in\partial D.$ $\Box$
Theorem 5.1.   Let $p\in(n-1,n],$ let $x_{0}\in\overline{\mathfrak{D}}$ and let $f_{m}\in\mathfrak{J}_{Q,\delta,p,E}(\mathfrak{D}),$ $m=1,2,\ldots,$ $f_{m}:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}$ be a sequence such that:
1) for every $m\in{\mathbb{N}},$ a domain $D_{m}$ is locally connected at $\partial D_{m};$
2) for every neighborhood $U$ of $x_{0}$ there is a neighborhood $V$ of $x_{0},$ $V\subset U,$ and $M=M(U,x_{0})\in{\mathbb{N}}$ such that $V\cap D_{m}$ is connected for every $m\geqslant M(U,x_{0});$
3) there is $\varepsilon_{0}>0$ such that $f^{\,-1}_{m}(K_{m})\subset{\mathbb{R}}^{n}\setminus B(x_{0},\varepsilon_{0}),$ where $K_{m}:=K_{f_{m}}$ is a continuum corresponding to the definition of the class $\mathfrak{J}_{Q,\delta,p,E}(\mathfrak{D})$ for $f_{m};$
4) the family $f_{m}(D_{m})$ is equi-uniform with respect to $p$-modulus over all $m\in{\mathbb{N}};$
5) at least one of two following conditions hold: $Q$ has a finite mean oscillation in $\overline{D_{0}},$ or relation (1.6) holds for every $x_{0}\in\overline{D_{0}}$ and some $\beta(x_{0})>0.$
Then the family $f_{m},$ $m=1,2,\ldots,$ is equicontinuous with respect to $x_{0},$ i.e., for any $\varepsilon>0$ there is $\delta=\delta(x_{0},\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(x^{\,\prime}))<\varepsilon$ whenever $x,x^{\,\prime}\in B(x_{0},\delta)\cap D_{m}$ and $m\in{\mathbb{N}}.$
[…]
Funding. The second author was supported by the National Research Foundation of Ukraine, Project number 2025.07/0014, Project name: ‘‘Modern problems of Mathematical Analysis and Geometric Function Theory’’.