### v3

# ON DISTORTION ESTIMATES UNDER MAPPINGS OF FAMILIES OF DOMAINS

Abstract
We consider families of mappings with moduli inequalities, having different definition domains. Under some additional assumptions we have proved that such families are uniformly equicontinuous. We have considered four main cases: when mappings are homeomorphisms and corresponding domains have simple geometry; when similar mappings have branch points; when domains with complex geometry are considered, but mappings still are homeomorphisms; and when similar mappings have branch points. Sequences of domains are generally assumed to converge to a kernel, and the characteristics of the mappings must satisfy certain conditions on their growth. In some of the four cases mentioned above, we also described properties of the limit mapping. We also obtained the correspondence of the boundary points of the kernel to the boundary points, and the inner points to the inner points.
2010 Mathematics Subject Classification: Primary 30C65; Secondary 31A15, 31B25
1 Introduction
This paper is devoted to the study of mappings satisfying upper distortion estimates of the modulus of families of paths. We consider the case when definition domains of mappings may be different. The main subject of this study is distortion estimates for the distance under mappings. For fixed domains, this problem has been studied by many authors, including both homeomorphisms and mappings with a branching. In particular, we could mentioned papers on this matter, see, e.g., [Cr1][Cr2], [MRSY], [M], [RS1] and [SevSkv]. Considering families of mappings of variable domains, we use some ideas from [Suv].
Let us recall some definitions. A Borel function $\rho:{\mathbb{R}}^{n}\,\rightarrow[0,\infty]$ is called admissible for the family $\Gamma$ of paths $\gamma$ in ${\mathbb{R}}^{n},$ if the relation
$\int\limits_{\gamma}\rho(x)\,|dx|\geqslant 1$
holds for all (locally rectifiable) paths $\gamma\in\Gamma.$ In this case, we write: $\rho\in{\rm adm}\,\Gamma.$ Let $p\geqslant 1,$ then $p$-modulus of $\Gamma$ is defined by the equality
$M_{p}(\Gamma)=\inf\limits_{\rho\in\,{\rm adm}\,\Gamma}\int\limits_{{\mathbb{R}}^{n}}\rho^{p}(x)\,dm(x)\,.$
We set $M(\Gamma):=M_{n}(\Gamma).$ Let $x_{0}\in{\mathbb{R}}^{n},$ $0<r_{1}<r_{2}<\infty,$
$S(x_{0},r)=\{x\,\in\,{\mathbb{R}}^{n}:|x-x_{0}|=r\}\,,\quad B(x_{0},r)=\{x\,\in\,{\mathbb{R}}^{n}:|x-x_{0}|<r\}$ (1.1)
and
$A=A(x_{0},r_{1},r_{2})=\left\{x\,\in\,{\mathbb{R}}^{n}:r_{1}<|x-x_{0}|<r_{2}\right\}\,.$ (1.2)
Given sets $E,$ $F\subset\overline{{\mathbb{R}}^{n}}$ and a domain $D\subset{\mathbb{R}}^{n}$ we denote by $\Gamma(E,F,D)$ a family of all paths $\gamma:[a,b]\rightarrow\overline{{\mathbb{R}}^{n}}$ such that $\gamma(a)\in E,\gamma(b)\in\,F$ and $\gamma(t)\in D$ for $t\in(a,b).$ Let $S_{i}=S(x_{0},r_{i}),$ $i=1,2,$ where spheres $S(x_{0},r_{i})$ centered at $x_{0}$ of the radius $r_{i}$ are defined in (1.1). Given $p\geqslant 1,$ a mapping $f:D\rightarrow\overline{{\mathbb{R}}^{n}}$ is called a ring $Q$-mapping at a point $x_{0}\in{\mathbb{R}}^{n}$ with respect to $p$-modulus, if the condition
$M_{p}(f(\Gamma(S_{1},S_{2},D)))\leqslant\int\limits_{A\cap D}Q(x)\cdot\eta^{p}(|x-x_{0}|)\,dm(x)$ (1.3)
holds for all $0<r_{1}<r_{2}<\infty$ and any Lebesgue measurable function $\eta:(r_{1},r_{2})\rightarrow[0,\infty]$ such that
$\int\limits_{r_{1}}^{r_{2}}\eta(r)\,dr\geqslant 1\,.$ (1.4)
The inequalities of the form (1.3) were established for many well-known classes of mappings, see e.g. [Va, Definition 13.1]), [MRSY, Theorems 8.1, 8.5] and [Sal, Corollary 2].
Recall that a mapping $f:D\rightarrow{\mathbb{R}}^{n}$ is called discrete if the pre-image $\{f^{-1}\left(y\right)\}$ of each point $y\,\in\,{\mathbb{R}}^{n}$ consists of isolated points, and is open if the image of any open set $U\subset D$ is an open set in ${\mathbb{R}}^{n}.$ A mapping $f:D\rightarrow{\mathbb{R}}^{n}$ is called closed if $f(A)$ is closed in $f(D)$ whenever $A$ is closed in $D.$ Later, in the extended space $\overline{{{\mathbb{R}}}^{n}}={{\mathbb{R}}}^{n}\cup\{\infty\}$ we use the spherical (chordal) metric $h(x,y)=|\pi(x)-\pi(y)|,$ where $\pi$ is a stereographic projection $\overline{{{\mathbb{R}}}^{n}}$ onto the sphere $S^{n}(\frac{1}{2}e_{n+1},\frac{1}{2})$ in ${{\mathbb{R}}}^{n+1},$ namely,
$h(x,\infty)=\frac{1}{\sqrt{1+{|x|}^{2}}}\,,\quad h(x,y)=\frac{|x-y|}{\sqrt{1+{|x|}^{2}}\sqrt{1+{|y|}^{2}}}\,,\quad x\neq\infty\neq y$
(see [Va, Definition 12.1]). In what follows, ${\rm Int\,}A$ denotes the set of inner points of the set $A\subset\overline{{\mathbb{R}}^{n}}.$ Recall that the set $U\subset\overline{{\mathbb{R}}^{n}}$ is neighborhood of the point $z_{0},$ if $z_{0}\in{\rm Int\,}A.$
Given $E_{1},E_{2}\subset\overline{{\mathbb{R}}^{n}}$ we set
$h(E_{1},E_{2})=\inf\limits_{x\in E_{1},y\in E_{2}}h(x,y)\,.$
Let $I$ be a fixed set of indices and let $D_{i},$ $i\in I,$ be some number of domains. Following [NP, Sect. 2.4], we say that a family of domains $\{D_{i}\}_{i\in I}$ is equi-uniform with respect to $p$-modulus if for any $r>0$ there exists a number $\delta>0$ such that the inequality $M_{p}(\Gamma(F^{\,*},F,D_{i}))\geqslant\delta$ holds for any $i\in I$ and any continua $F,F^{*}\subset D_{i}$ such that $h(F)\geqslant r$ and $h(F^{\,*})\geqslant r.$
Remark 1.1. Due to [MRSY, Section 3] we say that a boundary $D$ is called strongly accessible with respect to $p$-modulus at $x_{0}\in\partial D,$ if for any neighborhood $U$ of the point $x_{0}\in\partial D$ there is a neighborhood $V\subset U$ of this point, a compactum $F\subset D$ and a number $\delta>0$ such that $M_{p}(\Gamma(E,F,D))\geqslant\delta$ for any continua $E\subset D$ such that $E\cap\partial U\neq\varnothing\neq E\cap\partial V.$ The boundary of a domain $D$ is called strongly accessible with respect to $p$-modulus, if this is true for any $x_{0}\in\partial D.$ When $p=n,$ prefix ‘‘relative to $p$-modulus’’ is omitted. Uniform domains have strongly accessible boundaries, see e.g. [SevSkv, Remark 1]. The inverse statement is also true at least for $p=n,$ see [Sev2, Lemma 4.2].
Remark 1.2. Observe that, if $n=2,$ then every domain $D$ in ${\mathbb{R}}^{2}$ which has a finite number of boundary components is uniform if and only if $D$ is finitely connected on the boundary, see Corollary 6.8 in [Na2].
Set
$q_{x_{0}}(r):=\frac{1}{\omega_{n-1}r^{n-1}}\int\limits_{|x-x_{0}|=r}Q(x)\,dS\,$,
where $dS$ is the element of the surface area of $S,$ and
$q^{\,\prime}_{x_{0}}(r):=\frac{1}{\omega_{n-1}r^{n-1}}\int\limits_{|x-x_{0}|=r}Q^{\,\prime}(x)\,dS\,$,
$Q^{\,\prime}(x)=\max\{Q(x),1\}.$ A domain $D\subset{\mathbb{R}}^{n}$ is called locally connected at the point $x_{0}\in\partial D,$ if for any neighborhood $U$ of point $x_{0}$ there is a neighborhood $V\subset U$ of the same point such that $V\cap D$ is connected. A domain $D$ is called locally connected on $\partial D,$ if this domain is such at each point of its boundary.
Let $D_{m},$ $m=1,2,\ldots,$ be a sequence of domains in ${\mathbb{R}}^{n},$ containing a fixed point $A_{0}.$ If there exists a ball $B(A_{0},\rho),$ $\rho>0,$ belonging to all $D_{m},$ then the kernel of the sequence $D_{m},$ $m=1,2,\ldots,$ with respect to $A_{0}$ is the largest domain $D_{0}$ containing $x_{0}$ and such that for each compact set $E$ belonging to $D_{0}$ there is $N>0$ such that $E$ belongs to $D_{m}$ for all $m\geqslant N.$ A largest domain is one which contains any other domain having the same property. A sequence of domains $D_{m},$ $m=1,2,\ldots,$ converges to a kernel $D_{0}$ if any subsequence of $D_{m}$ has $D_{0}$ as its kernel.
Let $D_{m},$ $m=1,2,\ldots,$ be a sequence of domains which converges to a kernel $D_{0}.$ Then $D_{m}$ will be called regular with respect to $D_{0},$ if $D_{m}\subset D_{0}$ for all $m\in{\mathbb{N}}$ and, for every $x_{0}\in\overline{D_{0}}$ and for every neighborhood $U$ of $x_{0}$ there is a neighborhood $V$ of $x_{0},$ $V\subset U,$ and $M=M(U,x_{0})\in{\mathbb{N}}$ such that $V\cap D_{m}$ is a non-empty connected set for every $m\geqslant M(U,x_{0}).$
Given $p\geqslant 1,$ a number $\delta>0,$ a fixed domain $D_{0}\subset{\mathbb{R}}^{n},$ $n\geqslant 2,$ a sequence of domains $\mathfrak{D}=\{D_{m}\}_{m=1}^{\infty}$ the kernel of which is $D_{0},$ a non-degenerate continuum $A\subset D_{0},$ $A\subset\bigcap\limits_{m=1}^{\infty}D_{m},$ and a Lebesgue measurable function $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$, we denote by $\mathfrak{F}_{Q,A,p,\delta}(D_{0},\mathfrak{D})$ some a family of ring $Q$-homeomorphisms $f_{m}:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}$ satisfying (1.3)–(1.4) at any point $x_{0}\in\overline{D_{0}}$ for all $0<r_{1}<r_{2}<\infty$ such that $h(f_{m}(A))\geqslant\delta$ and $h(\overline{{\mathbb{R}}^{n}}\setminus f_{m}(D_{m}))\geqslant\delta.$ The concept of finite mean oscillation (FMO) used below can be found, for example, in [MRSY] or [Sev1]. The following statement is true.
Theorem 1.1.   Let $p\in(n-1,n]$ and let $f_{m}\in\mathfrak{F}_{Q,A,p,\delta}(D_{0},\mathfrak{D}),$ $m=1,2,\ldots,$ be a sequence such that:
1) the sequence of domains $D_{m}$ is regular with respect to $D_{0};$
2) for every $m\in{\mathbb{N}},$ a domain $D_{m}$ is locally connected on its boundary;
3) the family $f_{m}(D_{m})$ is equi-uniform with respect to $p$-modulus over all $m\in{\mathbb{N}};$
4) at least one of two following conditions hold: $Q$ has a finite mean oscillation in $\overline{D_{0}},$ or
$\int\limits_{0}^{\beta(x_{0})}\frac{dt}{t^{\frac{n-1}{p-1}}q_{x_{0}}^{\,\prime\,\frac{1}{p-1}}(t)}=\infty$ (1.5)
for every $x_{0}\in\overline{D_{0}}$ and some $\beta(x_{0})>0.$
I. Then the family $f_{m},$ $m=1,2,\ldots,$ is uniformly equicontinuous in $\mathfrak{D},$ i.e., for any $\varepsilon>0$ there is $\delta=\delta(\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(y))<\varepsilon$ whenever $x,y\in D_{m},$ $|x-y|<\delta$ and $m\in{\mathbb{N}}.$
II. Moreover, there is a subsequence $f_{m_{k}},$ $k=1,2,\dots,$ which converges to $f$ locally uniformly in $D_{0}.$ In this case, $f$ has a continuous boundary extension $f:\overline{D_{0}}\rightarrow\overline{{\mathbb{R}}^{n}}$ and $f$ is a homeomorphism in $D_{0}.$ If $A_{0}$ is some compactum in $D_{0},$ then there is $\delta_{2}>0$ and $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ for all $m\geqslant M_{0}$ and $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}.$
III. If $x_{m}\in D_{m},$ $m=1,2,\ldots,$ $f_{m}$ converges to $f$ locally uniformly in $D_{0}$ and $x_{m}\rightarrow x_{0}$ as $m\rightarrow\infty,$ then $f_{m}(x_{m})\rightarrow f(x_{0}).$ If $x_{0}\in\partial D_{0},$ then $f(x_{0})\in\partial f(D_{0}).$
We also prove a result similar to Theorem 1 for non-injective mappings.
Given $p\geqslant 1,$ a domain $D_{0}\subset{\mathbb{R}}^{n},$ a set $E\subset\overline{{\mathbb{R}}^{n}}$ and a number $\delta>0$, a fixed domain $D_{0}\subset{\mathbb{R}}^{n},$ $n\geqslant 2,$ a sequence of domains $\mathfrak{D}=\{D_{m}\}_{m=1}^{\infty}$ the kernel of which is $D_{0}$ and a Lebesgue measurable function $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$ we denote by $\mathfrak{R}_{Q,\delta,p,E}(D_{0},\mathfrak{D})$ the family of all open discrete closed mappings $f:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}\setminus E$ satisfying (1.3)–(1.4) for any $0<r_{1}<r_{2}<\infty$ at any point $x_{0}\in\overline{D_{0}}$ such that the following condition holds: for any domain $D^{\,\prime}_{m}:=f_{m}(D_{m})$ there exists a continuum $K_{m}\subset D^{\,\prime}_{m}$ such that $h(K_{m})\geqslant\delta$ and $h(f_{m}^{\,-1}(K_{m}),\partial D_{m})\geqslant\delta>0.$ The following statement holds.
Theorem 1.2.Let $p\in(n-1,n]$ and let $f_{m}\in\mathfrak{R}_{Q,\delta,p,E}(D_{0},\mathfrak{D}),$ $m=1,2,\ldots,$ be a sequence such that the conditions 1)–4) from Theorem 1 hold. Let $E$ be a set of a positive capacity for $p=n,$ and $E$ is arbitrary closed set for n1<p<n.n-1<p<n.
I. Then the family $f_{m},$ $m=1,2,\ldots,$ is uniformly equicontinuous in $\mathfrak{D},$ i.e., for any $\varepsilon>0$ there is $\delta=\delta(\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(y))<\varepsilon$ whenever $x,y\in D_{m},$ $|x-y|<\delta$ and $m\in{\mathbb{N}}.$ Moreover, there is a subsequence $f_{m_{k}},$ $k=1,2,\dots,$ which converges to $f$ locally uniformly in $D_{0}.$ In this case, $f$ has a continuous boundary extension $f:\overline{D_{0}}\rightarrow\overline{{\mathbb{R}}^{n}}.$ If $x_{m}\in D_{m},$ $m=1,2,\ldots,$ $f_{m}$ converges to $f$ locally uniformly in $D_{0}$ and $x_{m}\rightarrow x_{0}$ as $m\rightarrow\infty,$ then $f_{m}(x_{m})\rightarrow f(x_{0}).$
II. In addition, assume that
5) $p=n$ and there is $r>0,$ does not depending on $m\in{\mathbb{N}},$ such that $h(E)\geqslant r$ whenever $E$ is any component of $\partial f_{m}(D_{m}).$ If $A_{0}$ is some compactum in $D_{0},$ then there is $\delta_{2}>0$ and $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ and $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}.$ The mapping $f$ is boundary preserving: if $x_{0}\in\partial D_{0},$ then $f(x_{0})\in\partial f(D_{0}).$ If $\overline{B(x_{0},\varepsilon_{0})}\subset D_{0}\cap\bigcap\limits_{m=1}^{\infty}D_{m},$ then there is $\varepsilon_{1}>0$ does not depending on $m\in{\mathbb{N}}$ such that
$B_{h}(f_{m}(x_{0}),\varepsilon_{1})\subset f_{m}(D_{m})\,,\quad m=1,2,\ldots\,$, (1.6)
where $B_{h}(y_{0},r_{0})=\{y\in\overline{{\mathbb{R}}^{n}}:h(y,y_{0})<r_{0}\}.$
Results similar to Theorems 1 and 1 also hold for domains with prime ends. We now formulate these results.
Definitions related to prime ends of domains can be found in [Sev1] and are therefore omitted. Consider the following definition, which goes back to Näkki [Na1], cf. [KR]. The boundary of a domain $D$ in ${\mathbb{R}}^{n}$ is said to be locally quasiconformal if every $x_{0}\in\partial D$ has a neighborhood $U$ that admits a quasiconformal mapping $\varphi$ onto the unit ball ${\mathbb{B}}^{n}\subset{\mathbb{R}}^{n}$ such that $\varphi(\partial D\cap U)$ is the intersection of ${\mathbb{B}}^{n}$ and a coordinate hyperplane. The sequence of cuts $\sigma_{m},$ $m=1,2,\ldots,$ is called quasiconformally regular, if $\overline{\sigma_{m}}\cap\overline{\sigma_{m+1}}=\varnothing$ for $m\in{\mathbb{N}}$ and, in addition, $d(\sigma_{m})\rightarrow 0$ as $m\rightarrow\infty.$ If the end $K$ contains at least one quasiconformally regular chain, then $K$ will be called quasiconformally regular. We say that a bounded domain $D$ in ${\mathbb{R}}^{n}$ is quasiconformally regular, if $D$ can be quasiconformally mapped to a domain with a locally quasiconformal boundary whose closure is a compact in ${\mathbb{R}}^{n},$ and, besides that, every prime end in $D$ is quasiconformally regular. Note that space $\overline{D}_{P}=D\cup E_{D}$ is metric, which can be demonstrated as follows. If $g:D_{0}\rightarrow D$ is a quasiconformal mapping of a domain $D_{0}$ with a locally quasiconformal boundary onto some domain $D,$ then for $x,y\in\overline{D}_{P}$ we put:
$\rho(x,y):=|g^{\,-1}(x)-g^{\,-1}(y)|\,$, (1.7)
where the element $g^{\,-1}(x),$ $x\in E_{D},$ is to be understood as some (single) boundary point of $D_{0}.$ The specified boundary point is unique and well-defined by [IS, Theorem 2.1, Remark 2.1], cf. [Na1, Theorem 4.1]. It is easy to verify that $\rho$ in (1.7) is a metric on $\overline{D}_{P}.$ If $g_{*}$ is another quasiconformal mapping of a domain $D_{*}$ with locally quasiconformal boundary onto $D$, then the corresponding metric $\rho_{*}(p_{1},p_{2})=|{\widetilde{g_{*}}}^{-1}(p_{1})-{\widetilde{g_{*}}}^{-1}(p_{2})|$ generates the same convergence and, consequently, the same topology in $\overline{D}_{P}$ as $\rho_{0}$ because $g_{0}\circ g_{*}^{-1}$ is a quasiconformal mapping of $D_{*}$ and $D_{0}$, which extends, by Theorem 4.1 in [Na1], to a homeomorphism between $\overline{D_{*}}$ and $\overline{D_{0}}$. In the sequel, this topology in $\overline{D}_{P}$ will be called the topology of prime ends; the continuity of mappings $F\colon\overline{D}_{P}\rightarrow\overline{D^{\,\prime}}_{P}$ will be understood relative to this topology.
Let $D_{m},$ $m=1,2,\ldots,$ be a sequence of domains which converges to a kernel $D_{0}.$ Then $D_{m}$ will be called regular with respect to $D_{0}$ in the terms of prime ends, if $D_{m}\subset D_{0}$ for all $m\in{\mathbb{N}}$ and, for every $P_{0}\in E_{D_{0}}$ there is a sequence $\sigma_{k},$ $k=1,2,\ldots,$ with the following condition: if $d_{k}$ is a domain in $P_{0}$ then there is $M=M(k)$ such that $d_{k}\cap D_{m}$ is a non-empty connected set for every $m\geqslant M(k).$ The following result holds.
Theorem 1.3.   Let $p\in(n-1,n]$ and let $f_{m}\in\mathfrak{F}_{Q,A,p,\delta}(D_{0},\mathfrak{D}),$ $m=1,2,\ldots,$ be a sequence such that:
1) the sequence of domains $D_{m}$ is regular with respect to $D_{0}$ in terms of prime ends;
2) for every $m\in{\mathbb{N}},$ a domain $D_{m}$ is quasiconformally regular; $D_{0}$ is also quasiconformally regular;
3) the family $f_{m}(D_{m})$ is equi-uniform with respect to $p$-modulus over all $m\in{\mathbb{N}}$;
4) at least one of two following conditions hold: $Q$ has a finite mean oscillation in $\overline{D_{0}},$ or (1.5) holds for every $x_{0}\in\overline{D_{0}}$ and some $\beta(x_{0})>0.$
I. Then the family $f_{m},$ $m=1,2,\ldots,$ is uniformly equicontinuous in $\mathfrak{D}$ in terms of prime ends, i.e., for any $\varepsilon>0$ there is $\delta=\delta(\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(y))<\varepsilon$ whenever $x,y\in D_{m},$ $\rho(x,y)<\delta,$ $m\in{\mathbb{N}}$ and $\rho$ is one of the metric in $\overline{D_{0}}_{P}$ defined in (1.7).
II. Moreover, there is a subsequence $f_{m_{k}},$ $k=1,2,\dots,$ which converges to $f$ locally uniformly in $D_{0}$ with respect $\rho.$ In this case, $f$ has a continuous boundary extension $f:\overline{D_{0}}_{P}\rightarrow\overline{{\mathbb{R}}^{n}}$ and $f$ is a homeomorphism in $D_{0}.$ If $A_{0}$ is some compactum in $D_{0},$ then there is $\delta_{2}>0$ and $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ for all $m\geqslant M_{0}$ and $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}.$
III. If $x_{m}\in D_{m},$ $m=1,2,\ldots,$ $f_{m}$ converges to $f$ locally uniformly in $D_{0}$ and $x_{m}\rightarrow P_{0}$ as $m\rightarrow\infty,$ then $f_{m}(x_{m})\rightarrow f(P_{0}).$ If $P_{0}\in E_{D_{0}}:=\overline{D_{0}}_{P}\setminus D_{0},$ then $f(P_{0})\in\partial f(D_{0}).$
Theorem 1.4.Let $p\in(n-1,n]$ and let $f_{m}\in\mathfrak{R}_{Q,\delta,p,E}(D_{0},\mathfrak{D}),$ $m=1,2,\ldots,$ be a sequence such that the conditions 1)–4) from Theorem 1 hold. Let $E$ be a set of a positive capacity for $p=n,$ and $E$ is arbitrary closed set for n1<p<n.n-1<p<n.
I. Then the family $f_{m},$ $m=1,2,\ldots,$ is uniformly equicontinuous in $\mathfrak{D}$ with respect to prime ends, i.e., for any $\varepsilon>0$ there is $\delta=\delta(\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(y))<\varepsilon$ whenever $x,y\in D_{m},$ $\rho(x,y)<\delta,$ $m\in{\mathbb{N}}$ and $\rho$ is one of the metric in $\overline{D_{0}}_{P}.$ Moreover, there is a subsequence $f_{m_{k}},$ $k=1,2,\dots,$ which converges to $f$ locally uniformly in $D_{0}.$ In this case, $f$ has a continuous boundary extension $f:\overline{D_{0}}_{P}\rightarrow\overline{{\mathbb{R}}^{n}}.$ If $x_{m}\in D_{m},$ $m=1,2,\ldots,$ $f_{m}$ converges to $f$ locally uniformly in $D_{0}$ and $x_{m}\rightarrow P_{0}$ as $m\rightarrow\infty,$ then $f_{m}(x_{m})\rightarrow f(P_{0}).$
II. Besides that, if the condition 5) of Theorem 1 holds, then for any compactum $A_{0}\subset D_{0},$ $A_{0}\subset\bigcap\limits_{m=1}^{\infty}D_{m},$ there exists $\delta_{1}>0$ such that $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots.$ The mapping $f$ is boundary preserving: if $P_{0}\in E_{D_{0}},$ then $f(x_{0})\in\partial f(D_{0}).$ If $\overline{B(x_{0},\varepsilon_{0})}\subset D_{0}\cap\bigcap\limits_{m=1}^{\infty}D_{m},$ then there is $\varepsilon_{1}>0$ such that the relation (1.6) holds, where $B_{h}(y_{0},r_{0})=\{y\in\overline{{\mathbb{R}}^{n}}:h(y,y_{0})<r_{0}\}.$
2 Preliminaries
The following statement is true, see [Sev1, Lemma 2.1].
Lemma 2.1.  Let $p\geqslant 1,$ and let $f:D\rightarrow\overline{{\mathbb{R}}^{n}}$ be an open discrete ring $Q$-mapping at the point $b\in\partial D$ with respect to $p$-modulus, $b\neq\infty,$ $f(D)=D^{\,\prime}$, a domain $D$ is locally connected at the point $b,$ $C(f,\partial D)\subset\partial D^{\,\prime}$ and $D^{\,\prime}$ is strongly accessible with respect to the $p$-modulus at least at one point $y\in C(f,b),$ where $C(f,b)$ is a cluster set of $f$ at $b.$ Suppose that there is $\varepsilon_{0}>0$ and some positive measurable function $\psi:(0,\varepsilon_{0})\rightarrow(0,\infty)$ such that
$0<I(\varepsilon,\varepsilon_{0})=\int\limits_{\varepsilon}^{\varepsilon_{0}}\psi(t)\,dt<\infty$
for any $\varepsilon\in(0,\varepsilon_{0})$ and, in addition,
$\int\limits_{A(b,\varepsilon,\varepsilon_{0})}Q(x)\cdot\psi^{\,p}(|x-b|)\ dm(x)=o(I^{p}(\varepsilon,\varepsilon_{0}))\,$,
where $A:=A(b,\varepsilon,\varepsilon_{0})$ is defined in (1.2). Then $C(f,b)=\{y\}.$
Observe that, $C(f,\partial D)\subset\partial D^{\,\prime}$ for homeomorphisms of $D$ onto $D^{\,\prime}$, see e.g. [MRSY, Proposition 13.5].
We will prove the main assertions of the paper in a somewhat more general form. In this regard, consider the following definition. We will say that a point $x_{0}$ is a limit point of a family of domains $\mathfrak{D}$ in ${\mathbb{R}}^{n}$ if there exists a sequence $D_{m},$ $m=1,2,\ldots,$ in $\mathfrak{D}$ and a sequence $x_{m}\in D_{m},$ such that $x_{m}\rightarrow x_{0}$ as $m\rightarrow\infty.$ Let $\overline{\mathfrak{D}}$ be a set of all such points $x_{0}.$
Given $p\geqslant 1,$ a number $\delta>0,$ a family $\mathfrak{D}$ of domains in ${\mathbb{R}}^{n},$ $n\geqslant 2,$ and a Lebesgue measurable function $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$, we denote by $\mathfrak{H}^{p}_{Q,\delta}(\mathfrak{D})$ a family of all homeomorphisms $f:D\rightarrow\overline{{\mathbb{R}}^{n}}$ defined in some $D\in\mathfrak{D}$ and satisfying (1.3)–(1.4) at any point $x_{0}\in\overline{\mathfrak{D}},$ $0<r_{1}<r_{2}<\infty,$ for which there exists a continuum $A=A_{f}\subset D$ with $h(f(A))\geqslant\delta.$
The following lemma asserts the equicontinuity of a certain family of mappings at a fixed point. Unlike the main result of Theorem 1, this assertion does not assume the convergence of the sequence of domains to its kernel, and the corresponding domains in which the mappings are defined can be arbitrary. The following statement is true.
Lemma 2.2.   Let $p\in(n-1,n],$ let $x_{0}\in\overline{\mathfrak{D}}$ and let $f_{m}\in\mathfrak{H}^{p}_{Q,\delta}(\mathfrak{D}),$ $m=1,2,\ldots,$ $f_{m}:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}$ be a sequence such that:
1) for every $m\in{\mathbb{N}},$ $D_{m}$ is locally connected at $\partial D_{m};$
2) for every neighborhood $U$ of $x_{0}$ there is a neighborhood $V$ of $x_{0},$ $V\subset U,$ and $M=M(U,x_{0})\in{\mathbb{N}}$ such that $V\cap D_{m}$ is connected for every $m\geqslant M(U,x_{0});$
3) there is $\varepsilon_{0}>0$ such that $A_{m}:=A_{f_{m}}\subset{\mathbb{R}}^{n}\setminus B(x_{0},\varepsilon_{0}),$ where $A_{f_{m}}$ is a continuum corresponding to the definition of the class $\mathfrak{H}^{p}_{Q,\delta}(\mathfrak{D})$ for $f_{m};$
4) the family $f_{m}(D_{m})$ is equi-uniform with respect to $p$-modulus over all $m\in{\mathbb{N}};$
5) there exists $\varepsilon_{0}=\varepsilon_{0}(x_{0})>0$ and a Lebesgue measurable function $\psi:(0,\varepsilon_{0})\rightarrow[0,\infty]$ such that
$I(\varepsilon,\varepsilon_{0}):=\int\limits_{\varepsilon}^{\varepsilon_{0}}\psi(t)\,dt<\infty\quad\forall\,\,\varepsilon\in(0,\varepsilon_{0})\,,\quad I(\varepsilon,\varepsilon_{0})\rightarrow\infty\quad\text{as}\quad\varepsilon\rightarrow 0\,$, (2.1)
while
$\int\limits_{A(x_{0},\varepsilon,\varepsilon_{0})}Q(x)\cdot\psi^{\,p}(|x-x_{0}|)\,dm(x)=o(I^{p}(\varepsilon,\varepsilon_{0}))\,$, (2.2)
as $\varepsilon\rightarrow 0,$ where $A(x_{0},\varepsilon,\varepsilon_{0})$ is defined in (1.2). (If $x_{0}=\infty,$ (2.2) means the similar relation for $x_{0}=0$ with $\widetilde{Q}(x)=Q\left(\frac{x}{|x|^{2}}\right)$ instead of $Q(x)$).
Then the family $f_{m},$ $m=1,2,\ldots,$ is equicontinuous with respect to $x_{0},$ i.e., for any $\varepsilon>0$ there is $\delta=\delta(x_{0},\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(x^{\,\prime}))<\varepsilon$ whenever $x,x^{\,\prime}\in B(x_{0},\delta)\cap D_{m}$ and $m\in{\mathbb{N}}.$
Proof.   We will mainly use the methodology that has been used for mappings of a fixed domain, see, e.g., [SevSkv]. Assume the contrary. Then there exists $a>0$ such that for $\delta=1/k,$ $k=1,2,\ldots,$ there are $m_{k}\in{\mathbb{N}}$ and $x_{k},x^{\,\prime}_{k}\in B(x_{0},1/k)\cap D_{m_{k}}$ such that $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a.$ We may consider that the sequence $m_{k}$ is increasing by $k.$ Otherwise, $m_{k}=k_{0}$ for sufficiently large $k\in{\mathbb{N}}$ and some $k_{0}\in{\mathbb{N}},$ besides that, $x_{0}\in\overline{D_{k_{0}}}$ because $f_{k_{0}}$ is defined at $x_{k}\cap D_{k_{0}}$ and $x_{k}\rightarrow x_{0}$ as $k\rightarrow\infty.$ Now, from the assumption above we have that $h(f_{k_{0}}(x_{k}),f_{k_{0}}(x^{\,\prime}_{k}))\geqslant a.$ The latter contradicts with Lemma 2. So, $m_{k}$ may be chosen as an increasing sequence. Without loss of generality, going to renumbering, if required, we may consider that $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a$ holds for $m$ instead of $m_{k},$ i.e.,
$h(f_{m}(x_{m}),f_{m}(x^{\,\prime}_{m}))\geqslant a/2\qquad\forall\,\,m\in{\mathbb{N}}\,.$ (2.3)
We may assume that $x_{0}\neq\infty.$ Now, we use the condition 2): for every neighborhood $U$ of $x_{0}$ there is a neighborhood $V$ of $x_{0},$ $V\subset U,$ and $M=M(U,x_{0})\in{\mathbb{N}}$ such that $V\cap D_{m}$ is connected for every $m\geqslant M(U,x_{0}).$ Thus, given a ball $U_{1}=B(x_{0},2^{\,-1})$ there is $V_{1}\subset U_{1}$ and $M(1)\in{\mathbb{N}}$ such that $D_{m}\cap V_{1}$ is connected for $m\geqslant M(1).$ Since $x_{m},x_{m}^{\,\prime}\rightarrow x_{0}$ as $m\rightarrow\infty,$ there is $m_{1}\geqslant M(1)$ such that $x_{m_{1}},x^{\,\prime}_{m_{1}}\in V_{1}\cap D_{m_{1}}.$ Follow, given a ball $U_{2}=B(x_{0},2^{\,-2})$ there is $V_{2}\subset U_{2}$ and $M(2)\in{\mathbb{N}}$ such that $D_{m}\cap V_{2}$ is connected for $m\geqslant M(2).$ Since $x_{m},x_{m}^{\,\prime}\rightarrow x_{0}$ as $m\rightarrow\infty,$ there is $m_{2}\geqslant\max\{M(2),m_{1}\}$ such that $x_{m_{2}},x^{\,\prime}_{m_{2}}\in V_{2}\cap D_{m_{2}}.$ Etc. In general, given a ball $U_{k}=B(x_{0},2^{\,-k})$ there is $V_{k}\subset U_{k}$ and $M(k)\in{\mathbb{N}}$ such that $D_{m}\cap V_{k}$ is connected for $m\geqslant M(k).$ Since $x_{m},x_{m}^{\,\prime}\rightarrow x_{0}$ as $m\rightarrow\infty,$ there is $m_{k}\geqslant\max\{M(k),m_{k-1}\}$ such that $x_{m_{k}},x^{\,\prime}_{m_{k}}\in V_{k}\cap D_{m_{k}}.$ Relabeling, if required, we may consider that the same sequences $x_{m}$ and $x^{\prime}_{m}$ satisfy the above conditions, i.e., given a ball $U_{m}=B(x_{0},2^{\,-m}),$ $m=1,2,\ldots,$ there is $V_{m}\subset U_{m}$ such that $x_{m},x^{\,\prime}_{m}\in D_{m}\cap V_{m}$ while $D_{m}\cap V_{m}$ is connected. We join the points $x_{m}$ and $x^{\,\prime}_{m}$ by a path $\gamma_{m}:[0,1]\rightarrow D_{m}\cap V_{m}$ such that $\gamma_{m}(0)=x_{m},$ $\gamma_{m}(1)=x^{\,\prime}_{m}$ and $\gamma_{m}(t)\in D_{m}\cap V_{m}$ for $t\in(0,1),$ see Figure 1.
Figure 1: To prove Lemma 2
Denote by $C_{m}$ the image of the path $\gamma_{m}(t)$ under $f_{m}.$ It follows from relation (2.3) that
$h(C_{m})\geqslant a/2\qquad\forall\,m\in{\mathbb{N}}\,$, (2.4)
where $h(C_{m})$ denotes the chordal diameter of the set $C_{m}.$
Let $\Gamma_{m}$ be a family of paths joining $|\gamma_{m}|$ and $A_{m}$ in $D_{m},$ where $A_{m}$ is a continuum from the condition 3). By the definition of the class $\mathfrak{H}^{p}_{Q,\delta}(\mathfrak{D}),$ the relations (1.3)–(1.4) hold at $x_{0}.$ Therefore,
$M_{p}(f_{m}(\Gamma_{m}))\leqslant\int\limits_{A(x_{0},2^{\,-m},\varepsilon_{0})}Q(x)\cdot\eta^{p}(|x-x_{0}|)\ dm(x)$ (2.5)
for every Lebesgue measurable function $\eta:(2^{\,-m},\varepsilon_{0})\rightarrow[0,\infty]$ such that $\int\limits_{2^{\,-m}}^{\varepsilon_{0}}\eta(r)\,dr\geqslant 1.$ Observe that, the function
[…]
where $I(\varepsilon,\varepsilon_{0}):=\int\limits_{\varepsilon}^{\varepsilon_{0}}\psi(t)\,dt,$ satisfies the condition (1.4) with $r_{1}:=2^{\,-m},$ $r_{2}:=\varepsilon_{0}.$ Therefore, it follows from (2.2) and (2.5) that
$M_{p}(f_{m}(\Gamma_{m}))\leqslant\alpha(2^{\,-m})\rightarrow 0$ (2.6)
as $m\rightarrow\infty,$ where $\alpha(\varepsilon)$ is some nonnegative function tending to zero as $\varepsilon\rightarrow 0,$ which exists due to the condition (2.2).
On the other hand, observe that $f_{m}(\Gamma_{m})=\Gamma(C_{m},f_{m}(A_{m}),D_{m}^{\,\prime}).$ By the hypothesis of the lemma, $h(f_{m}(A_{m}))\geqslant\delta$ for all $m\in{\mathbb{N}}.$ Therefore, due to (2.4), $h(f_{m}(A_{m}))\geqslant\delta_{1}$ and $h(C_{m})\geqslant\delta_{1},$ where $\delta_{1}:=\min\{\delta,a/2\}.$ Using the fact that the domains $D_{m}^{\,\prime}:=f_{m}(D_{m})$ are equi-uniform with respect to $p$-modulus, we conclude that there exists $\sigma>0$ such that
$M_{p}(f_{m}(\Gamma_{m}))=M_{p}(\Gamma(C_{m},f_{m}(A_{m}),D_{m}^{\,\prime}))\geqslant\sigma\qquad\forall\,\,m\in{\mathbb{N}}\,$,
which contradicts the condition (2.6). The obtained contradiction indicates that the assumption of the absence of equicontinuity of the family $\mathfrak{H}^{p}_{Q,\delta}(\mathfrak{D})$ at $x_{0}$ was incorrect. The obtained contradiction completes the proof of the lemma. $\Box$
The following statement may be found in [Sev1, Lemmas 1.3, 1.4].
Proposition 2.1.   Let $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty],$ $n\geqslant 2,$ be a Lebesgue measurable function and let $x_{0}\in{\mathbb{R}}^{n}.$ Let $0<p\leqslant n.$ Assume that either of the following conditions holds
(a) $Q\in FMO(x_{0}),$
(b) $q_{x_{0}}(r)\,=\,O\left(\left[\log{\frac{1}{r}}\right]^{n-1}\right)$ as $r\rightarrow 0,$
(c) for some small $\delta_{0}=\delta_{0}(x_{0})>0$ we have the relations
$\int\limits_{\delta}^{\delta_{0}}\frac{dt}{t^{\frac{n-1}{p-1}}q_{x_{0}}^{\frac{1}{p-1}}(t)}<\infty,\qquad 0<\delta<\delta_{0},\qquad\int\limits_{0}^{\delta_{0}}\frac{dt}{t^{\frac{n-1}{p-1}}q_{x_{0}}^{\frac{1}{p-1}}(t)}=\infty\,.$
Then there exist a number $\varepsilon_{0}\in(0,1)$ and a function $\psi\geqslant 0$ satisfying (2.1) such that the relation (2.2) holds.
Uniting Lemma 2 and Proposition 2, we obtain the following.
Theorem 2.1.   Let $p\in(n-1,n],$ let $x_{0}\in\overline{\mathfrak{D}}$ and let $f_{m}\in\mathfrak{H}^{p}_{Q,\delta}(\mathfrak{D}),$ $m=1,2,\ldots,$ $f_{m}:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}$ be a sequence such that:
1) for every $m\in{\mathbb{N}},$ $D_{m}$ is locally connected at $\partial D_{m};$
2) for every neighborhood $U$ of $x_{0}$ there is a neighborhood $V$ of $x_{0},$ $V\subset U,$ and $M=M(U,x_{0})\in{\mathbb{N}}$ such that $V\cap D_{m}$ is connected for every $m\geqslant M=M(U,x_{0});$
3) there is $\varepsilon_{0}>0$ such that $A_{m}:=A_{f_{m}}\subset{\mathbb{R}}^{n}\setminus B(x_{0},\varepsilon_{0}),$ where $A_{f_{m}}$ is a continuum corresponding to the definition of the class $\mathfrak{H}^{p}_{Q,\delta}(\mathfrak{D})$ for $f_{m};$
4) the family $f_{m}(D_{m})$ is equi-uniform with respect to $p$-modulus over all $m\in{\mathbb{N}};$
5) at least one of two following conditions hold: $Q$ has a finite mean oscillation in $\overline{D_{0}},$ or (1.5) holds for every $x_{0}\in\overline{D_{0}}$ and some $\beta(x_{0})>0.$
Then the family $f_{m},$ $m=1,2,\ldots,$ is equicontinuous with respect to $x_{0},$ i.e., for any $\varepsilon>0$ there is $\delta=\delta(x_{0},\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(x^{\,\prime}))<\varepsilon$ whenever $x,x^{\,\prime}\in B(x_{0},\delta)\cap D_{m}$ and $m\in{\mathbb{N}}.$
3 On image of the continuum under open discrete and closed mappings
The following statement holds, see, e.g., [Ku, Theorem 1.I.5.46]).
Proposition 3.1.   Let $A$ be a set in a topological space $X.$ If the set $C$ is connected and $C\cap A\neq\varnothing\neq C\setminus A,$ then $C\cap\partial A\neq\varnothing.$
Let $D$ be a domain in ${\mathbb{R}}^{n},$ $n\geqslant 2,$ and let $f:D\rightarrow{\mathbb{R}}^{n}$ (or $f:D\rightarrow\overline{{\mathbb{R}}^{n}})$ be a discrete open mapping, $\beta:[a,\,b)\rightarrow{\mathbb{R}}^{n}$ be a path, and $x\in\,f^{-1}(\beta(a)).$ A path $\alpha:[a,\,b)\rightarrow D$ is called a total $f$-lifting of $\beta$ starting at $x,$ if $(1)\quad\alpha(a)=x\,;$ $(2)\quad(f\circ\alpha)(t)=\beta(t)$ for any $t\in[a,b).$
Proposition 3.2.   Let $f:D\rightarrow{\mathbb{R}}^{n}$ be a discrete open and closed (boundary preserving) mapping, $\beta:[a,\,b)\rightarrow f(D)$ be a path, and $x\in\,f^{\,-1}\left(\beta(a)\right).$ Then $\beta$ has a total $f$-lifting $\alpha:[a,b)\rightarrow D$ starting at $x$ (see, e.g., [Vu, Lemma 3.7]). Moreover, if $\beta(t)\rightarrow\partial f(D)$ as $t\rightarrow b,$ then $\alpha(t)\rightarrow\partial D$ as $t\rightarrow b$ (see, e.g., Lemma 3.12 in [MRV]).
The following statement holds, see e.g. [MRSY, Theorem 7.2].
Lemma 3.1. (Väisälä’s lemma on the weak flatness of inner points).   Let $n\geqslant 2$, let $D$ be a domain in $\overline{{\mathbb{R}}^{n}},$ and let $x_{0}\in D.$ Then for each $P>0$ and each neighborhood $U$ of point $x_{0}$ there is a neighborhood $V\subset U$ of the same point such that $M(\Gamma(E,F,D))>P$ for any continua $E,F\subset D$ intersecting $\partial U$ and $\partial V.$
[…]
Lemma 3.2.  Let $D_{0}\subset{\mathbb{R}}^{n},$ be a domain in ${\mathbb{R}}^{n},$ $n\geqslant 2,$ and let $\mathfrak{D}=\{D_{m}\}_{m=1}^{\infty}$ be a sequence of domains the kernel of which is $D_{0}.$ Let $A$ be a non-degenerate continuum in $D_{0},$ $A\subset\bigcap\limits_{m=1}^{\infty}D_{m},$ and let $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$ be a Lebesgue measurable function. In addition, let $f_{m}:D_{m}\rightarrow{\mathbb{R}}^{n}$ be a sequence of open discrete and closed mappings satisfying relations (1.3)–(1.4) at any point $x_{0}\in\overline{D_{0}}$ for all $0<r_{1}<r_{2}<d(x_{0},\partial D_{m})$ such that $h(f_{m}(A))\geqslant\delta>0$ for any $m=1,2,\ldots$ and $p=n$ for some $\delta>0.$ Assume that, $Q$ has a finite mean oscillation in $A,$ or the relation (1.5) holds for every $x_{0}\in A$ and some $\beta(x_{0})>0.$ If there is $r>0$ such that $h(E)\geqslant r$ whenever $E$ is any component of $\partial f_{m}(D_{m}),$ then there exists $\delta_{1}>0$ such that $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots.$
Proof.   Observe that, $h(A,\partial D_{m})\geqslant r_{*}>0$ for some $r_{*}>0,$ and all $m\in{\mathbb{N}}$ except finite numbers. Otherwise, there are sequences $x_{m}\in A$ and $y_{m}\in\partial D_{m}$ with $h(x_{m},y_{m})\rightarrow 0$ as $m\rightarrow\infty.$ We may consider that $x_{m},y_{m}\rightarrow x_{0}$ as $m\rightarrow\infty.$ Obviously, $x_{0}\in A$ by the closeness of $A.$ Now, there is $\varepsilon_{0}>0$ such that $\overline{B(x_{0},\varepsilon_{0})}\subset D_{0}$ and $\overline{B(x_{0},\varepsilon_{0})}$ belongs to all $D_{m}$ except of finite numbers. However, the latter contradicts the condition $y_{m}\rightarrow x_{0}$ as $m\rightarrow\infty$ while $y_{m}\in\partial D_{m},$ because in this case $\overline{B(x_{0},\varepsilon_{0})}\cap\partial D_{m}\neq\varnothing$ for infinitely many $m\in{\mathbb{N}}.$
We firstly consider the case when the relation (1.5) holds. Since for any $0<\varepsilon_{1}<\widetilde{\varepsilon_{1}}<\min\{\beta(x_{0}),\frac{1}{2}r_{*}\}$ we have that
$\int\limits_{\varepsilon_{1}}^{\widetilde{\varepsilon_{1}}}\frac{dt}{tq_{x_{0}}^{\,\prime\,\frac{1}{n-1}}(t)}\leqslant\int\limits_{\varepsilon_{1}}^{\widetilde{\varepsilon_{1}}}\frac{dt}{t}<\infty\,$,
the relation (1.5) implies that, given $0<\widetilde{\varepsilon_{1}}<r_{*}\leqslant h(A,\partial D_{m})$ we may find $\varepsilon_{1}=\varepsilon_{1}(x_{0})$ such that $\int\limits_{\varepsilon_{1}}^{\widetilde{\varepsilon_{1}}}\frac{dt}{tq_{x_{0}}^{\,\prime\,\frac{1}{n-1}}(t)}>1\,.$ We cover $A$ with the balls $B(x_{0},\varepsilon_{1}),$ $x_{0}\in A,$ where $\varepsilon_{1}$ as above. By Heine-Borel-Lebesgue lemma we may find a finite subcover $B(x_{1},\varepsilon_{1}),$ $B(x_{2},\varepsilon_{2}),\ldots,B(x_{N},\varepsilon_{N})$ such that $A\subset\bigcup\limits_{i=1}^{n}B(x_{i},\varepsilon_{i})\,$, $A\subset\bigcup\limits_{i=1}^{n}B(x_{i},2\varepsilon_{i})\,.$ By the choice of $\varepsilon_{i}>0,$ $\varepsilon_{i}<\min\{\beta(x_{0}),\frac{1}{2}r_{*}\},$ we have that $B(x_{i},2\varepsilon_{i})\subset D_{m}.$ Since $D_{m}$ is a domain in ${\mathbb{R}}^{n},$ $m=1,2,\ldots,$ $\partial D_{m}\neq\varnothing.$ In addition, since $C(\partial D_{m},f_{m})\subset\partial f_{m}(D_{m})$ whenever $f_{m}$ is open, discrete and closed (see [Vu, Theorem 3.3]), $\partial f_{m}(D_{m})\neq\varnothing,$ as well. Thus, the quantity $h(f_{m}(A),\partial f_{m}(D_{m}))$ is well-defined, so the formulation of the lemma is correct.
Let us prove Lemma 3 by the contradiction, partially using the approach of the proof of Lemma 4.1 in [SevSkv]. Suppose that the conclusion of the lemma is not true. Then for each $k\in{\mathbb{N}}$ there is some number $m=m_{k}$ such that $h(f_{m_{k}}(A),\partial f_{m_{k}}(D_{m_{k}}))<1/k.$ In order not to complicate the notation, we will further assume that $h(f_{m}(A),\partial f_{m}(D_{m}))<1/m,$ $m=1,2,\ldots.$ Note that the set $f_{m}(A)$ is compact as a continuous image of a compact set $A\subset D_{m}$ under the mapping $f_{m}.$ In this case, there are elements $x_{m}\in f_{m}(A)$ and $y_{m}\in\partial f_{m}(D_{m})$ such that $h(f_{m}(A),\partial f_{m}(D_{m}))=h(x_{m},y_{m})<1/m.$ Let $E_{m}$ be a component of $\partial f_{m}(D_{m})$ containing $y_{m}.$ Due to the compactness of $\overline{{\mathbb{R}}^{n}},$ we may assume that $y_{m}\rightarrow y_{0}$ as $m\rightarrow\infty;$ then also $x_{m}\rightarrow y_{0}$ as $m\rightarrow\infty.$
Put $P>0$ and $U=B_{h}(y_{0},r_{0})=\{y\in\overline{{\mathbb{R}}^{n}}:h(y,y_{0})<r_{0}\},$ where $2r_{0}:=\min\{r/2,\delta/2\},$ where $r$ and $\delta$ are numbers from the condition of the lemma. Observe that $E_{m}\cap U\neq\varnothing\neq E_{m}\setminus U$ for sufficiently large $m\in{\mathbb{N}},$ since $y_{m}\rightarrow y_{0}$ as $m\rightarrow\infty,$ $y_{m}\in E_{m};$ besides that, $h(E_{m})\geqslant r>r/2\geqslant 2r_{0}$ and $h(U)\leqslant 2r_{0}.$ Observe that, $E_{m}$ is closed (see [Ku, Theorem 1.III.5.46]). Now, $E_{m}$ is a continuum, so that $E_{m}\cap\partial U\neq\varnothing$ by Proposition 3. Similarly, $f_{m}(A)\cap U\neq\varnothing\neq f_{m}(A)\setminus U$ for sufficiently large $m\in{\mathbb{N}},$ since $x_{m}\rightarrow y_{0}$ as $m\rightarrow\infty,$ $x_{m}\in f_{m}(A);$ besides that $h(f_{m}(A))\geqslant\delta>\delta/2\geqslant 2r_{0}$ and $h(U)\leqslant 2r_{0}.$ Since $f_{m}(A)$ is a continuum, $f_{m}(A)\cap\partial U\neq\varnothing$ by Proposition 3. By the proving above, $E_{m}\cap\partial U\neq\varnothing\neq f_{m}(A)\cap\partial U$ for sufficiently large $m\in{\mathbb{N}}.$ By Lemma 3 there is $\subset U, VV$ is a neighborhood of $y_{0},$ such that
$M(\Gamma(E,F,\overline{{\mathbb{R}}^{n}}))>P$ (3.1)
for any continua $E,F\subset\overline{{\mathbb{R}}^{n}}$ with $E\cap\partial U\neq\varnothing\neq E\cap\partial V$ and $F\cap\partial U\neq\varnothing\neq F\cap\partial V.$ Arguing similarly to above, we may prove that $E_{m}\cap\partial V\neq\varnothing\neq f_{m}(A)\cap\partial V$ for sufficiently large $m\in{\mathbb{N}}.$ Thus, by (3.1)
$M(\Gamma(f_{m}(A),E_{m},\overline{{\mathbb{R}}^{n}}))>P$ (3.2)
for sufficiently large $m=1,2,\ldots.$ We now prove that the relation (3.2) contradicts the definition of $f_{m}$ in (1.3)–(1.4).
Indeed, let $\gamma:[0,1]\rightarrow\overline{{\mathbb{R}}^{n}}$ be a path in $\Gamma(f_{m}(A),E_{m},\overline{{\mathbb{R}}^{n}}),$ i.e., $\gamma(0)\in f_{m}(A),$ $\gamma(1)\in E_{m}$ and $\gamma(t)\in\overline{{\mathbb{R}}^{n}}$ for $t\in(0,1).$ Let $t_{m}=\sup\limits_{\gamma(t)\in f_{m}(D_{m})}t$ and let $\alpha_{m}(t)=\gamma|_{[0,t_{m})}.$ Let $\Gamma_{m}$ consists of all such paths $\alpha_{m},$ now $\Gamma(f_{m}(A),E_{m},\overline{{\mathbb{R}}^{n}})>\Gamma_{m}$ and by the minorization principle of the modulus (see [Fu, Theorem 1])
$M(\Gamma_{m})\geqslant M(\Gamma(f_{m}(A),E_{m},\overline{{\mathbb{R}}^{n}}))\,.$ (3.3)
Let $\Delta_{m}$ be a family of all total $f_{m}$-liftings of $\Gamma_{m}$ starting at $A$ (it exists by Proposition 3). Now
$M(f_{m}(\Delta_{m}))=M(\Gamma_{m})\,.$ (3.4)
Let $\beta_{m}:[0,t_{m})\rightarrow D_{m},$ $\beta_{m}\in\Delta_{m},$ $f_{m}\circ\beta_{m}=\alpha_{m}.$ By Proposition 3 $\beta_{m}(t)\rightarrow\partial D_{m}$ as $t\rightarrow t_{m}.$
We now show that
$\Delta_{m}>\bigcup\limits_{i=1}^{N}\Gamma(S(x_{i},\varepsilon_{i}),S(x_{i},2\varepsilon_{i}),A(x_{i},\varepsilon_{i},2\varepsilon_{i}))\,.$ (3.5)
Indeed, let $\gamma\in\Delta_{m},$ in other words, $\gamma:[0,t_{m})\rightarrow D_{m},$ $\gamma(0)\in A$ and $\gamma(t)\rightarrow\partial D_{m}$ as $t\rightarrow t_{m}-0.$ Now, there is $1\leqslant i\leqslant N$ such that $|\gamma|\cap B(x_{i},\varepsilon_{i})\neq\varnothing\neq|\gamma|\cap(D_{m}\setminus B(x_{i},\varepsilon_{i})).$ Therefore, by Proposition 3 there is $0<t^{\,*}_{1}<t_{m}$ such that $\gamma(t^{\,*}_{1})\in S(x_{i},\varepsilon_{i}).$ We may assume that $\gamma(t)\not\in B(x_{i},\varepsilon_{i})$ for $t>t^{\,*}_{1}.$ Put $\gamma_{1}:=\gamma|_{[t^{\,*}_{1},t_{m}]}.$ Similarly, $|\gamma_{1}|\cap B(x_{i},2\varepsilon_{i})\neq\varnothing\neq|\gamma_{1}|\cap(D_{m}\setminus B(x_{i},2\varepsilon_{i})).$ By Proposition 3 there is $t^{\,*}_{1}<t^{\,*}_{2}<t_{m}$ with $\gamma(t^{\,*}_{2})\in S(x_{i},2\varepsilon_{i}).$ We may assume that $\gamma(t)\in B(x_{i},2\varepsilon_{i})$ for $t<t^{\,*}_{2}.$ Put $\gamma_{2}:=\gamma|_{[t^{\,*}_{1},t^{\,*}_{2}]}.$ Then, the path $\gamma_{2}$ is a subpath of $\gamma,$ which belongs to the family $\Gamma(S(x_{i},\varepsilon_{i}),S(x_{i},2\varepsilon_{i}),A(x_{i},\varepsilon_{i},2\varepsilon_{i})).$ Thus, the relation (3.5) is established.
On the other hand, we set
$\psi_{i}(t)=\left\{\begin{array}[]{rr}1/[tq_{x_{i}}^{\,\prime\frac{1}{n-1}}(t)]\ ,&\ t\in(\varepsilon_{i},2\varepsilon_{i})\ ,\\ 0\ ,&\ t\notin(\varepsilon_{i},2\varepsilon_{i})\ ,\end{array}\right$.
where $q^{\,\prime}_{x_{i}}(r):=\frac{1}{\omega_{n-1}r^{n-1}}\int\limits_{|x-x_{i}|=r}Q^{\,\prime}(x)\,dS,$ $Q^{\,\prime}(x)=\max\{Q(x),1\}.$ Now,
$\int\limits_{A(x_{i},\varepsilon_{i},2\varepsilon_{i})}Q^{\,\prime}(x)\cdot\psi^{n}(|x-x_{i}|)\,dm(x)=\omega_{n-1}I_{i}\,$, (3.6)
where $I_{i}=\int\limits_{\varepsilon_{i}}^{2\varepsilon_{i}}\frac{dt}{tq_{x_{i}}^{{}^{\,\prime}\frac{1}{n-1}}(t)}.$ Note that, the function $\eta_{i}(t)=\psi(t)/I_{i},$ $t\in(\varepsilon_{i},2\varepsilon_{i}),$ satisfies (1.4), because $\int\limits_{\varepsilon_{i}}^{2\varepsilon_{i}}\eta_{i}(t)\,dt=1.$ Now, it follows from (3.6) that
$M(f_{m}(\Gamma(S(x_{i},\varepsilon_{i}),S(x_{i},2\varepsilon_{i}),A(x_{i},\varepsilon_{i},2\varepsilon_{i})))\leqslant\frac{\omega_{n-1}}{I_{i}^{n-1}}<\infty\,$, (3.7)
where $I_{i}=I_{i}(x_{i},\varepsilon_{i},2\varepsilon_{i})=\int\limits_{\varepsilon_{i}}^{2\varepsilon_{i}}\frac{dr}{rq_{x_{i}}^{\,\prime\frac{1}{n-1}}(r)}$ and $I_{i}>0$ by the construction.
Finally, by (3.3), (3.4), (3.5) and (3.7) we obtain that
$M(\Gamma(f_{m}(A),E_{m},\overline{{\mathbb{R}}^{n}}))\leqslant M(\Gamma_{m})$
$=M(f_{m}(\Delta_{m}))\leqslant\sum\limits_{i=1}^{N}M(f_{m}(\Gamma(S(x_{i},\varepsilon_{i}),S(x_{i},2\varepsilon_{i}),A(x_{i},\varepsilon_{i},2\varepsilon_{i}))))$ (3.8)
$\leqslant\sum\limits_{i=1}^{N}\frac{\omega_{n-1}}{I_{i}^{n-1}}:=C<\infty\,.$
Since $P$ in (3.2) may be done arbitrary large, the relations (3.2) and (3.8) contradict each other. This completes the proof for the case when the relation (1.5) holds. Observe that, the case when $Q$ has a finite mean oscillation in $A$ is a particular case of the relation (1.5) (see the last part of the proof of Theorem 5.5 in [Sev1]). The proof is wholly complete. $\Box$
Corollary 3.1.   Assume that, under conditions of Lemma 3, $A_{0}$ is some another compactum in $D_{0}.$ Then there is $\delta_{2}>0$ and $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ for all $m\geqslant M_{0}$ and $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}.$
Proof.   Since $D_{m}$ converges to $D_{0}$ as its kernel, there is $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ for all $m\geqslant M_{0}.$ It remains to show that $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}$ and some $\delta_{2}>0.$ Assume the contrary, i.e., $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\rightarrow 0$ as $m\rightarrow\infty.$ Now, $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))=h(x_{m},y_{m})\rightarrow 0$ as $m\rightarrow\infty$ for $x_{m}\in A_{0}$ and $y_{m}\in f_{m}(D_{m}).$ We may consider that $x_{m}\rightarrow x_{0}$ as $m\rightarrow\infty,$ $x_{0}\in A_{0}.$ Put $\varepsilon_{0}>0$ such that $\overline{B(x_{0},\varepsilon_{0})}\subset D_{0}.$ We join $x_{0}$ with $A$ by a path $\gamma$ in $D_{0}.$ Now, $B:=A\cup|\gamma|\cup\overline{B(x_{0},\varepsilon_{0})}$ is a continuum in $D_{0}$ such that $B\subset D_{m}$ for sufficiently large $m,$ $x_{m}\in B$ for sufficiently large $m$ and $h(f_{m}(B))\geqslant h(f_{m}(B))\geqslant\delta$ for above $m.$ By Lemma 3, $h(f_{m}(B),\partial f_{m}(D_{m}))\geqslant\delta_{*}>0$ for $m\geqslant m_{3},$ $m_{3}\in{\mathbb{N}}.$ However, the latter contradicts with $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))=h(x_{m},y_{m})\rightarrow 0$ as $m\rightarrow\infty,$ because obviously $h(f_{m}(B),\partial f_{m}(D_{m}))\leqslant h(x_{m},y_{m})\rightarrow 0$ as $m\rightarrow\infty.$ The obtained contradiction completes the proof. $\Box$
4 Proof of Theorem 1
Proposition 4.1.   Let $D$ and $G$ be domains in ${\mathbb{R}}^{n},$ $n\geqslant 2,$ such that $D\subset G,$ and let $K$ be a compactum in $D.$ Then $h(K,\partial D)\leqslant h(K,\partial G).$
Proof.   Since $\overline{{\mathbb{R}}^{n}}$ is connected, $\partial D\neq\varnothing.$ By the same reason, $\partial G\neq\varnothing.$ Thus, $h(K,\partial G)$ and $h(K,\partial D)$ are well-defined.
Let now $h(K,\partial D)=h(x_{0},y_{0}),$ $x_{0}\in K$ and $y_{0}\in\partial D,$ and let $h(K,\partial G)=h(z_{0},z_{0}),$ $z_{0}\in K$ and $w_{0}\in\partial G.$ We need to prove that $h(x_{0},y_{0})\leqslant h(z_{0},w_{0}).$ We prove this by the contradiction: assume that $h(z_{0},w_{0})<h(x_{0},y_{0}).$ Consider the ball $\overline{B_{h}(z_{0},h(z_{0},w_{0}))}.$ Since $D\subset G,$ the point $w_{0}$ cannot be inner point of $D.$ Let us join the points $z_{0}$ and $w_{0}$ by a path $\gamma$ in $\overline{B_{h}(z_{0},h(z_{0},w_{0}))}.$ Now, $|\gamma|\cap D\neq\varnothing\neq|\gamma|\setminus D.$ By Proposition 3 there is $v_{0}\in\partial D\cap|\gamma|.$ Thus, $v_{0}\in\overline{B_{h}(z_{0},h(z_{0},w_{0}))}$ that means $h(z_{0},v_{0})\leqslant h(z_{0},w_{0}).$ Since $z_{0}\in K$ and $v_{0}\in\partial D,$ the latter means that
$h(K,\partial D)\leqslant h(z_{0},v_{0})\leqslant h(z_{0},w_{0})<h(x_{0},y_{0})=h(K,\partial D)\,.$
The obtained contradiction completes the proof. $\Box$
Proof of Theorem 1. I. Let us prove the item I of Theorem 1 by the contradiction. Assume that the assertion of Theorem 1 is false. Then for some $\varepsilon_{0}>0$ and for numbers $\delta_{k}=\frac{1}{k},$ $k=1,2,\ldots,$ there are $m_{k}\in{\mathbb{N}}$ and $x_{k},y_{k}\in D_{m_{k}}$ such that $h(f_{m_{k}}(x_{k}),f_{m_{k}}(y_{k}))\geqslant\varepsilon_{0}$ whenever $|x_{k}-y_{k}|<\frac{1}{k}.$ By Lemma 2 every $f_{m_{k}}$ has a continuous extension to $\partial D_{m_{k}},$ so that $f_{m_{k}}$ is uniformly continuous in $\overline{D_{m_{k}}}.$ Thus, we may consider that the sequence $m_{k},$ $k=1,2,\ldots,$ is increasing. Without loss of generality, we may consider that the above holds for $m=1,2,\ldots,$ but not for $m_{k}.$ In particular,
$h(f_{m}(x_{m}),f_{m}(y_{m}))\geqslant\varepsilon_{0}\,,\qquad m\in{\mathbb{N}}\,.$ (4.1)
By the compactness of $\overline{{\mathbb{R}}^{n}},$ we may consider that $x_{m},y_{m}\rightarrow x_{0}$ as $m\rightarrow\infty.$ There are two cases: 1) $x_{0}\in D_{0},$ 2) $x_{0}\in\partial D_{0}.$ In the first case, when $x_{0}$ is the inner point of $D_{0},$ all mappings $f_{m}$ except of finite numbers are defined in some ball $B(x_{0},\varepsilon_{0})\subset D_{0}\cap D_{m},$ $m\geqslant M_{0}.$ Now, $f_{m}$ is equicontinuous at $x_{0}$ (see [RS1, Theorems 6.1, 6.5] for $p=n$ and [Sev1, Theorems 4.1, 4.3] for n1<p<n.n-1<p<n. Now, by the triangle inequality,
$h(f_{m}(x_{m}),f_{m}(y_{m}))\leqslant h(f_{m}(x_{m}),f_{m}(x_{0}))+h(f_{m}(x_{0}),f_{m}(y_{m}))\,.$ (4.2)
In this case, the right side in (4.2) tends to zero as $m\rightarrow\infty$ which contradicts with (4.1), and the proof finishes.
Let us consider the case 2) $x_{0}\in\partial D_{0}.$ Observe that $A_{m}:=A\subset{\mathbb{R}}^{n}\setminus B(x_{0},\varepsilon_{0})$ for some $\varepsilon_{0}>0,$ because the $A$ is a continuum in $D_{0}.$ Thus, all of conditions of Theorem 2 are satisfied. By this theorem, for any $\varepsilon>0$ there is $\delta=\delta(x_{0},\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(x^{\,\prime}))<\varepsilon$ whenever $x,x^{\,\prime}\in B(x_{0},\delta)\cap D_{m}$ and $m\in{\mathbb{N}}.$ In particular, it follows from this theorem that there is $M=M(\varepsilon)\in{\mathbb{N}}$ such that
$h(f_{m}(x_{m}),f_{m}(y_{m}))<\varepsilon\,,\qquad m\geqslant M(\varepsilon)\,.$ (4.3)
The relations (4.1) and (4.3) contradict each other and prove the uniform equicontinuity of $f_{m}.$
II. Given a compactum $K$ in $D_{0},$ there is $m_{0}=m_{0}(K)$ such that $K\subset D_{m}$ for $m\geqslant m_{0}(K).$ Now, by the proving above, $f_{m},$ $m=1,2,\ldots,$ is uniformly equicontinuous family on $K.$ By Arzela-Ascoli Theorem (see. e.g., [Va, Theorem 20.4]) $f_{m}$ forms a normal family on $K,$ i.e., there exists a subsequence $f_{m_{k}},$ $k=1,2,\ldots,$ such that $f_{m_{k}}(x)\rightarrow f(x)$ as $k\rightarrow\infty.$ Since $K$ is an arbitrary compactum in $D_{0},$ the mapping $f(x)$ is defined in all $D_{0}.$ Now, we prove that $f$ has a continuous extension at every point $x_{0}\in\partial D_{0}.$ By the proving above, given $\varepsilon>0$ there is $\delta=\delta(\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(y))<\varepsilon$ whenever $x,y\in D_{m},$ $|x-y|<\delta$ and $m\in{\mathbb{N}}.$ Fix $x_{0}\in\partial D_{0}.$ Let $\delta_{*}=\delta_{*}(\varepsilon)=\delta/2$ and let $x,y\in B(x_{0},\delta_{*})\cap D_{0}.$ Now, by the triangle inequality $|x-y|<2\delta_{*}=\delta.$ Since $x,y\in D_{0},$ there is $m_{0}=m_{0}(x,y)\in{\mathbb{N}}$ such that $x,y\in D_{0}$ for every $m\geqslant m_{0}.$ Taking into account the mentioned above, we obtain that $h(f_{m}(x),f_{m}(y))<\varepsilon,$ $m\geqslant m_{0}.$ Taking the limit as $m\rightarrow\infty$ in the latter inequality, we obtain that $h(f(x),f(y))\leqslant\varepsilon.$ Now, $f$ satisfies the Cauchy condition at $x_{0},$ so that $f$ has a limit at $x_{0},$ as required.
Let $A_{0}$ be a compactum in $D_{0}$ and let $G$ be an arbitrary domain in $D_{0}$ with $A_{0}\cup A\overline{G}\subset D_{0}.$ Since $D_{m}$ converges to $D_{0}$ as its kernel, there is $m_{0}\in{\mathbb{N}}$ such that $G\cup A_{0}\subset D_{m},$ $m\geqslant m_{0}.$ Due to the above, $f_{m_{k}}$ converges to $f$ uniformly in $G.$ Now, $f$ is either a homeomorphism, or a constant in $G$ (see [RS2, Theorems 4.1 and 4.2] for $p=n$ and [Cr2, Theorem 1] for n1<p<nn-1<p<n). Due to the condition $h(f_{m}(A))\geqslant\delta,$ $f$ is homeomorphism in $G.$ Since $G$ is arbitrary compact subdomain in $D_{0},$ $f$ is continuous and injective in $D_{0}.$ Now, $f$ is a homeomorphism in $D_{0},$ as required (see, e.g., [RS2, Corollary 3.1]).
Let us to prove that, there is $\delta_{2}>0$ and $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ for all $m\geqslant M_{0}$ and $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}.$ Without loss of generality, we may consider that $f_{m}$ converges to $f$ locally uniformly as $m\rightarrow\infty.$ Let $\varepsilon:=\frac{1}{2}h(f(A_{0}),\partial f(G))$ ($\varepsilon>0$ because $f$ is a homeomorphism in $G$ and, consequently, $f(A_{0})\in\textrm{Int}\,f(G)$). By the triangle inequality and due to the uniform convergence of $f_{m}$ to $f,$ given $\varepsilon>0$ there is $M=M(\varepsilon)$ such that $h(f_{m}(x),f(x))<\varepsilon/2$ for any $x\in A_{0}.$ Let $y\in\partial f(G).$ By the triangle inequality
$h(f_{m}(x),y)\geqslant h(y,f(x))-h(f(x),f_{m}(x))\geqslant\varepsilon-\varepsilon/2=\varepsilon/2=\frac{1}{4}h(f(A_{0}),\partial f(G))\,.$
Thus, taking the $\inf$ over $x\in A_{0}$ and $y\in\partial f(G),$ we obtain that $h(f_{m}(A_{0}),\partial f(G))\geqslant\frac{1}{4}h(f(A_{0}),\partial f(G)):=\delta_{*}.$ Let now
$h(f_{m}(A_{0}),\partial f_{m}(G))=h(f_{m}(A_{0}),f_{m}(\partial G))=h(f_{m}(x_{m}),f_{m}(x^{\,\prime}_{m}))\,$,
where $x_{m}\in A_{0}$ and $x^{\,\prime}_{m}\in\partial G.$ Now, by the triangle inequality and by the locally uniform convergence of $f_{m}$ to $f,$ we obtain that
$h(f_{m}(x_{m}),f_{m}(x^{\,\prime}_{m}))\geqslant h(f_{m}(x_{m}),f(x^{\,\prime}_{m}))-h(f(x^{\,\prime}_{m}),f_{m}(x^{\,\prime}_{m}))\geqslant\delta_{*}-\delta_{*}/2=\delta_{*}/2>0$
for sufficiently large $m\in{\mathbb{N}}.$ The latter inequality implies that $h(f_{m}(A_{0}),\partial f_{m}(G))\geqslant\delta_{*}/2:=\delta_{2}.$ Finally, since $G\subset D_{m}$ for sufficiently large $m\in{\mathbb{N}},$ by Proposition 4 we obtain that $h(f_{m}(A_{0}),\partial f_{m}(G))\geqslant\delta_{2}$ for above $m,$ as required.
III. Let $x_{m}\in D_{m},$ $m=1,2,\ldots,$ $x_{m}\rightarrow x_{0}$ and $f_{m}\rightarrow f$ locally uniformly in $D_{0}$ as $m\rightarrow\infty.$ We need to show that $f_{m}(x_{m})\rightarrow f(x_{0}).$ If $x_{0}\in D_{0},$ the relation $f_{m}(x_{m})\rightarrow f(x_{0})$ as $m\rightarrow\infty$ directly follows from the locally uniform convergence of $f_{m}$ to $f$ in $D_{0}.$ Now, let $x_{0}\in\partial D_{0}.$ We lead the proof by the contradiction, i.e., assume that $h(f_{m_{k}}(x_{m_{k}}),f(x_{0}))\geqslant\varepsilon_{0}>0$ for some an increasing sequence of numbers $m_{k}=1,2,\ldots,$ and some $\varepsilon_{0}>0.$ Again, going to renumbering, if required, we may assume that
$h(f_{m}(x_{m}),f(x_{0}))\geqslant\varepsilon_{0}>0\,,\qquad m=1,2,\ldots\,.$ (4.4)
Consider some a sequence $y_{m}\rightarrow x_{0},$ $y_{m}\in D_{0},$ $m\rightarrow\infty.$ Now, given $m\in{\mathbb{N}},$ by the locally uniform convergence of $f_{m}$ to $f,$ there is $k=k_{m}$ such that $h(f_{k_{m}}(y_{k_{m}}),f(y_{k_{m}}))<1/m\,.$ We may consider that $k_{m}$ is an increasing sequence by $m=1,2,\ldots.$ Let $z_{m}:=y_{k_{m}}.$ Now, $z_{m}\rightarrow x_{0}$ and $f(z_{m})\rightarrow f(z_{0})$ as $m\rightarrow\infty$ because $f$ is continuous at $x_{0}$ (see item II). Now, by the triangle inequality and I we obtain that
$h(f_{k_{m}}(x_{k_{m}}),f(x_{0}))\leqslant h(f_{k_{m}}(x_{k_{m}}),f_{k_{m}}(z_{m}))$
$+h(f_{k_{m}}(z_{m}),f(z_{m}))+h(f(z_{m}),f(x_{0}))\rightarrow 0$ (4.5)
as $m\rightarrow\infty.$ The relation (4.5) contradicts with (4.4), that proves that $f_{m}(x_{m})\rightarrow f(x_{0})$ as $m\rightarrow\infty.$ Finally, $f(x_{0})\in\partial f(D_{0})$ whenever $x_{0}\in\partial D_{0},$ because $f$ is a homeomorphism in $D_{0}$ (see, e.g., [MRSY, Proposition 13.5]). Theorem is completely proved. $\Box$
5 Main Lemma for mappings with a branching
Given $p\geqslant 1,$ a number $\delta>0,$ a family $\mathfrak{D}$ of domains in ${\mathbb{R}}^{n},$ $n\geqslant 2,$ and a Lebesgue measurable function $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$, we denote by $\mathfrak{J}_{Q,\delta,p}(\mathfrak{D})$ the family of all open discrete closed ring mappings $f:D_{f}\rightarrow\overline{{\mathbb{R}}^{n}}$ defined in some a domain $D_{f}\in\mathfrak{D}$ and satisfying (1.3)–(1.4) at any point $x_{0}\in\overline{\mathfrak{D}}$ such that the following condition holds: for any domain $D^{\,\prime}_{f}:=f(D_{f})$ there exists a continuum $K_{f}\subset D^{\,\prime}_{f}$ such that $h(K_{f})\geqslant\delta$ and $h(f^{\,-1}(K_{f}),\partial D_{f})\geqslant\delta>0.$ The following lemma is an analog of Lemma 2 for mappings which are not homeomorphisms.
Lemma 5.1.   Let $p\in(n-1,n],$ let $x_{0}\in\overline{\mathfrak{D}}$ and let $f_{m}\in\mathfrak{J}_{Q,\delta,p}(\mathfrak{D}),$ $m=1,2,\ldots,$ $f_{m}:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}$ be a sequence such that:
1) for every $m\in{\mathbb{N}},$ a domain $D_{m}$ is locally connected at $\partial D_{m};$
2) for every neighborhood $U$ of $x_{0}$ there is a neighborhood $V$ of $x_{0},$ $V\subset U,$ and $M=M(U,x_{0})\in{\mathbb{N}}$ such that $V\cap D_{m}$ is connected for every $m\geqslant M(U,x_{0});$
3) there is $\varepsilon_{0}>0$ such that $f^{\,-1}_{m}(K_{m})\subset{\mathbb{R}}^{n}\setminus B(x_{0},\varepsilon_{0}),$ where $K_{m}:=K_{f_{m}}$ is a continuum corresponding to the definition of the class $\mathfrak{J}_{Q,\delta,p}(\mathfrak{D})$ for $f_{m};$
4) the family $f_{m}(D_{m})$ is equi-uniform with respect to $p$-modulus over all $m\in{\mathbb{N}};$
5) there exists $\varepsilon_{0}=\varepsilon_{0}(x_{0})>0$ and a Lebesgue measurable function $\psi:(0,\varepsilon_{0})\rightarrow[0,\infty]$ such that relations (2.1)–(2.2) hold as $\varepsilon\rightarrow 0,$ where $A(x_{0},\varepsilon,\varepsilon_{0})$ is defined in (1.2).
Then the family $f_{m},$ $m=1,2,\ldots,$ is equicontinuous with respect to $x_{0},$ i.e., for any $\varepsilon>0$ there is $\delta=\delta(x_{0},\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(x^{\,\prime}))<\varepsilon$ whenever $x,x^{\,\prime}\in B(x_{0},\delta)\cap D_{m}$ and $m\in{\mathbb{N}}.$
Proof.   Suppose the contrary. Then there exists $x_{0}\in\overline{\mathfrak{D}}$ and a number $a>0$ with the following property: for $\delta=1/k,$ $k=1,2,\ldots,$ there are $m_{k}\in{\mathbb{N}}$ and $x_{k},x^{\,\prime}_{k}\in B(x_{0},1/k)\cap D_{m_{k}}$ such that $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a.$ Arguing similarly to the proof of Lemma 2, we may consider that the sequence $m_{k}$ is increasing by $k.$ Without loss of generality, going to renumbering, if required, we may consider that $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a$ holds for $m$ instead of $m_{k},$ i.e., (2.3) is satisfied. We may assume that $x_{0}\neq\infty.$ Now, we use the condition 2): for every neighborhood $U$ of $x_{0}$ there is a neighborhood $V$ of $x_{0},$ $V\subset U,$ and $M=M(k)$ such that $V\cap D_{m}$ is connected for every $m\geqslant M(k).$ Thus, given a ball $U_{1}=B(x_{0},2^{\,-1})$ there is $V_{1}\subset U_{1}$ and $M(1)\in{\mathbb{N}}$ such that $D_{m}\cap V_{1}$ is connected for $m\geqslant M(1).$ Since $x_{m},x_{m}^{\,\prime}\rightarrow x_{0}$ as $m\rightarrow\infty,$ there is $m_{1}\geqslant M(1)$ such that $x_{m_{1}},x^{\,\prime}_{m_{1}}\in V_{1}\cap D_{m_{1}}.$ Follow, given a ball $U_{2}=B(x_{0},2^{\,-2})$ there is $V_{2}\subset U_{2}$ and $M(2)\in{\mathbb{N}}$ such that $D_{m}\cap V_{2}$ is connected for $m\geqslant M(2).$ Since $x_{m},x_{m}^{\,\prime}\rightarrow x_{0}$ as $m\rightarrow\infty,$ there is $m_{2}\geqslant\max\{M(2),m_{1}\}$ such that $x_{m_{2}},x^{\,\prime}_{m_{2}}\in V_{2}\cap D_{m_{2}}.$ Etc. In general, given a ball $U_{k}=B(x_{0},2^{\,-k})$ there is $V_{k}\subset U_{k}$ and $M(k)\in{\mathbb{N}}$ such that $D_{m}\cap V_{k}$ is connected for $m\geqslant M(k).$ Since $x_{m},x_{m}^{\,\prime}\rightarrow x_{0}$ as $m\rightarrow\infty,$ there is $m_{k}\geqslant\max\{M(k),m_{k-1}\}$ such that $x_{m_{k}},x^{\,\prime}_{m_{k}}\in V_{k}\cap D_{m_{k}}.$ Relabeling, if required, we may consider that the same sequences $x_{m}$ and $x^{\prime}_{m}$ satisfy the above conditions, i.e., given a ball $U_{m}=B(x_{0},2^{\,-m}),$ $m=1,2,\ldots,$ there is $V_{m}\subset U_{m}$ such that $x_{m},x^{\,\prime}_{m}\in D_{m}\cap V_{m}$ while $D_{m}\cap V_{m}$ is connected. We join the points $x_{m}$ and $x^{\,\prime}_{m}$ by a path $\gamma_{m}:[0,1]\rightarrow D_{m}\cap V_{m}$ such that $\gamma_{m}(0)=x_{m},$ $\gamma_{m}(1)=x^{\,\prime}_{m}$ and $\gamma_{m}(t)\in D_{m}\cap V_{m}$ for $t\in(0,1),$ see Figure 4.
Figure 2: To prove Lemma 5
Denote by $C_{m}$ the image of the path $\gamma_{m}$ under the mapping $f_{m}.$ It follows from relation (2.3) that a condition (2.4) is satisfied, where $h(C_{m})$ denotes the chordal diameter of the set $C_{m}.$
By the definition of the family $\mathfrak{J}_{Q,\delta,p}(\mathfrak{D}),$ for any $f_{m}\in\mathfrak{J}_{Q,\delta,p}(\mathfrak{D})$ and any domain $D^{\,\prime}_{m}:=f_{m}(D_{m})$ there exists a continuum $K_{m}\subset D^{\,\prime}_{m}$ such that $h(K_{m})\geqslant\delta$ and $h(f_{m}^{\,-1}(K_{m}),\partial D_{m})\geqslant\delta>0.$ Since by the condition of the lemma the domains $D^{\,\prime}_{m}$ form the equi-uniform family with respect to $p$-modulus and due to (2.4), we obtain that the inequality
$M_{p}(\Gamma(K_{m},C_{m},D^{\,\prime}_{m}))\geqslant b$ (5.1)
holds for all $m=1,2,\ldots$ and some $b>0.$ Consider the family $\Gamma_{m},$ consisting of all paths $\beta:[0,1)\rightarrow D^{\,\prime}_{m},$ where $\beta(0)\in C_{m}$ and $\beta(t)\rightarrow p\in K_{m}$ as $t\rightarrow 1.$ Let $\Gamma^{*}_{m}$ be a family of all total $f_{m}$-liftings $\alpha:[0,1)\rightarrow D$ of the family $\Gamma_{m}$ starting at $|\gamma_{m}|$ (see Proposition 3). Since $f_{m}$ is closed, we have: $\alpha(t)\rightarrow f^{\,-1}_{m}(K_{m}),$ where $f^{\,-1}_{m}(K_{m})$ is the total preimage of the continuum $K_{m}$ under the mapping $f_{m}.$ By (1.3), it follows that
$M_{p}(f_{m}(\Gamma_{m}^{*}))$
$\leqslant M_{p}(f_{m}(\Gamma(|\gamma_{m}|,f^{\,-1}_{m}(K_{m}),D_{m})))\leqslant\int\limits_{A(x_{0},2^{\,-m},\varepsilon_{0})}Q(x)\cdot\eta^{p}(|x-x_{0}|)\ dm(x)$ (5.2)
for each measurable function $\eta:(2^{\,-m},\varepsilon_{0})\rightarrow[0,\infty],$ such that $\int\limits_{2^{\,-m}}^{\varepsilon_{0}}\eta(r)\,dr\geqslant 1.$ Note that, the function
[…]
where $I(\varepsilon,\varepsilon_{0}):=\int\limits_{\varepsilon}^{\varepsilon_{0}}\psi(t)\,dt,$ satisfies the normalization condition (1.4). Therefore from the conditions (2.1)–(2.2) and (5.2) it follows that
$M_{p}(f_{m}(\Gamma^{\,*}_{m}))\leqslant\alpha(2^{\,-m})\rightarrow 0$ (5.3)
as $m\rightarrow\infty,$ where $\alpha(\varepsilon)$ is some nonnegative function tending to zero as $\varepsilon\rightarrow 0,$ which exists by the condition (2.2). In addition, $f_{m}(\Gamma^{\,*}_{m})=\Gamma_{m}$ and $M_{p}(\Gamma_{m})=M_{p}(\Gamma(K_{m},C_{m},D^{\,\prime}_{m})).$ Thus
$M_{p}(f_{m}(\Gamma^{\,*}_{m}))=M_{p}(\Gamma(K_{m},C_{m},D^{\,\prime}_{m}))\,.$ (5.4)
However, the relation (5.3) contradicts with (5.1) and (5.4). The resulting contradiction indicates that the assumption (2.3) was incorrect, and, therefore, the family $\mathfrak{J}_{Q,\delta,p}(\mathfrak{D})$ is equicontinuous at $x_{0}\in\partial D.$ $\Box$
Theorem 5.1.   Let $p\in(n-1,n],$ let $x_{0}\in\overline{\mathfrak{D}}$ and let $f_{m}\in\mathfrak{J}_{Q,\delta,p}(\mathfrak{D}),$ $m=1,2,\ldots,$ $f_{m}:D_{m}\rightarrow\overline{{\mathbb{R}}^{n}}$ be a sequence such that:
1) for every $m\in{\mathbb{N}},$ a domain $D_{m}$ is locally connected at $\partial D_{m};$
2) for every neighborhood $U$ of $x_{0}$ there is a neighborhood $V$ of $x_{0},$ $V\subset U,$ and $M=M(U,x_{0})\in{\mathbb{N}}$ such that $V\cap D_{m}$ is connected for every $m\geqslant M(U,x_{0});$
3) there is $\varepsilon_{0}>0$ such that $f^{\,-1}_{m}(K_{m})\subset{\mathbb{R}}^{n}\setminus B(x_{0},\varepsilon_{0}),$ where $K_{m}:=K_{f_{m}}$ is a continuum corresponding to the definition of the class $\mathfrak{J}_{Q,\delta,p}(\mathfrak{D})$ for $f_{m};$
4) the family $f_{m}(D_{m})$ is equi-uniform with respect to $p$-modulus over all $m\in{\mathbb{N}};$
5) at least one of two following conditions hold: $Q$ has a finite mean oscillation in $\overline{D_{0}},$ or relation (1.5) holds for every $x_{0}\in\overline{D_{0}}$ and some $\beta(x_{0})>0.$
Then the family $f_{m},$ $m=1,2,\ldots,$ is equicontinuous with respect to $x_{0},$ i.e., for any $\varepsilon>0$ there is $\delta=\delta(x_{0},\varepsilon)>0$ such that $h(f_{m}(x),f_{m}(x^{\,\prime}))<\varepsilon$ whenever $x,x^{\,\prime}\in B(x_{0},\delta)\cap D_{m}$ and $m\in{\mathbb{N}}.$
[…]
Proof of Theorem 1. The proofs of the statements in item I are based on Theorem 5 (instead of Theorem 2) and are completely similar to the corresponding proofs of Theorem 1.
II. Let $A_{0}$ be a compactum in $D_{0}.$ We show that there is $\delta_{2}>0$ and $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ and $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}.$ Note that, the relation $A_{0}\subset D_{m}$ for sufficiently large $m\in{\mathbb{N}}$ holds obviously, because $D_{m}$ converges to $D_{0}$ as its kernel by the assumption. Now, we show that $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for sufficiently large $m\in{\mathbb{N}}.$ Assume the contrary. Then $h(f_{m_{k}}(A_{0}),\partial f_{m_{k}}(D_{m_{k}}))\rightarrow 0$ as $k\rightarrow\infty$ for some increasing a sequence of numbers $m_{k},$ $k=1,2,\ldots.$ Going to renumbering, if required, we may consider that the latter holds for $m$ instead of $m_{k},$ i.e.,
$h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\rightarrow 0\,,\quad m\rightarrow\infty\,.$ (5.5)
Let $K_{m}\subset D^{\,\prime}_{m}$ be a continuum from the definition of the class $\mathfrak{R}_{Q,\delta,p,E}(D_{0},\mathfrak{D})$ such that $h(K_{m})\geqslant\delta$ and $h(f_{m}^{\,-1}(K_{m}),\partial D_{m})\geqslant\delta>0.$ Let $h(K_{m})=h(y_{m},z_{m}),$ where $y_{m},z_{m}\in K_{m}.$ Now, $y_{m}=f_{m}(x_{m})$ and $z_{m}=f_{m}(x^{\,\prime}_{m})$ for some $x_{m},x^{\,\prime}_{m}\in f_{m}^{\,-1}(K_{m}).$ We may consider that $x_{m}\rightarrow x_{0}$ and $x^{\,\prime}_{m}\rightarrow x^{\,\prime}_{0}$ as $m\rightarrow\infty.$ Due to the condition $h(f_{m}^{\,-1}(K_{m}),\partial D_{m})\geqslant\delta>0$ we have that $h(x_{m},\partial D_{m})\geqslant\delta$ and $h(x^{\,\prime}_{m},\partial D_{m})\geqslant\delta,$ $m=1,2,\ldots.$ Since by the assumption $D_{m}\subset D_{0},$ $m=1,2,\ldots,$ by Proposition 4 $h(x_{m},\partial D_{0})\geqslant\delta$ and $h(x^{\,\prime}_{m},\partial D_{0})\geqslant\delta,$ $m=1,2,\ldots.$ Now, $x_{0},x^{\,\prime}_{0}\in D_{0}.$ Let $\sigma_{1},\sigma_{2}>0$ be numbers such that $\overline{B(x_{0},\sigma_{1})}\subset D_{0}$ and $\overline{B(x^{\,\prime}_{0},\sigma_{2})}\subset D_{0}.$ We may consider that $x_{m}\in\overline{B(x_{0},\sigma_{1})}$ and $x^{\,\prime}_{m}\in\overline{B(x^{\,\prime}_{0},\sigma_{2})}$ for any $m\in{\mathbb{N}}.$ Now, we join the points $x_{0}$ and $x^{\,\prime}_{0}$ by a path $\gamma$ in $D_{0}.$ Let $A:=\overline{B(x_{0},\sigma_{1})}\cup|\gamma|\cup\overline{B(x^{\,\prime}_{0},\sigma_{2})},$ now $A$ is a continuum in $D_{0}.$ Since $D_{m}$ converges to $D_{0}$ as its kernel, we may consider that $A\subset D_{m}$ for any $m=1,2,\ldots.$ Observe that, $h(f_{m}(A))\geqslant h(f_{m}(x_{m}),f_{m}(x^{\,\prime}_{m}))=h(y_{m},z_{m})\geqslant\delta>0.$ Now, by Lemma 3 $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{*}>0$ for all $m\in{\mathbb{N}}$ and some $\delta_{*}>0.$ Besides that, by Corollary 3 $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{**}>0$ for all $m\in{\mathbb{N}}$ and some $\delta_{**}>0.$ The latter contradicts with the assumption (5.5). Thus, $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for sufficiently large $m\in{\mathbb{N}},$ as required.
Let us to prove that, the limit mapping $f$ is boundary preserving. In other words, wee need to prove that, if $x_{0}\in\partial D_{0},$ then $f(x_{0})\in\partial f(D_{0}).$ Assume the contrary: let $x_{0}\in\partial D_{0},$ however, $f(x_{0})\in{\rm Int\,}f(D_{0}).$ Now, $\partial D_{m}\cap U\neq\varnothing$ for any neighborhood $U$ of $x_{0}$ and for infinitely many $m=1,2,\ldots.$ So, we may construct a sequence $x_{k}\in\partial D_{m_{k}},$ $k=1,2,\ldots,$ such that $x_{k}\rightarrow x_{0}$ as $k\rightarrow\infty.$ For simplicity, we may consider that $x_{m}\in\partial D_{m}$ and $x_{m}\rightarrow x_{0}$ as $m\rightarrow\infty.$ Since $f_{m}$ are closed, they are boundary preserving (see, e.g., [Vu, Theorem 3.3]). Thus, for any $m\in{\mathbb{N}}$ we may find $y_{m}\in D_{m}$ such that $|x_{m}-y_{m}|<\frac{1}{m}$ and $h(f_{m}(y_{m}),\partial f_{m}(D_{m}))<\frac{1}{m}.$ Let $h(f_{m}(y_{m}),\partial f_{m}(D_{m}))=h(f_{m}(y_{m}),z_{m}).$ Let $A_{0}:=\{y_{0}\}$ and $y_{0}$ is some point of $D_{0}$ with $f(y_{0})=f(x_{0}).$ Now, $y_{0}\in D_{m}$ for all $m\in{\mathbb{N}}$ except a finite number. By the triangle inequality and proving above we obtain that
$h(f_{m}(y_{0}),z_{m})\leqslant h(f_{m}(y_{0}),f(x_{0}))+h(f(x_{0}),f_{m}(y_{m}))+h(f_{m}(y_{m}),z_{m})\rightarrow 0$
as $m\rightarrow\infty.$ The latter contradicts with the statement that $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0},$
Let us to prove the relation (1.6). We lead the proof by the contradiction, i.e., assume that (1.6) does not hold. Now, given $k\in{\mathbb{N}}$ there is $m=m_{k}\in{\mathbb{N}}$ such that $B_{h}(f_{m_{k}}(x_{0}),1/k)\setminus f_{m_{k}}(D_{m_{k}})\neq\varnothing$ whenever $\overline{B(x_{0},\varepsilon_{0})}\subset D_{0}\cap\bigcap\limits_{m=1}^{\infty}D_{m}.$ Let $p_{k}\in B_{h}(f_{m_{k}}(x_{0}),1/k)\setminus f_{m_{k}}(D_{m_{k}}).$ We join the points $f_{m_{k}}(x_{0})$ and $p_{k}$ inside the ball $B_{h}(f_{m_{k}}(x_{0}),1/k)$ by a path $\gamma_{k}.$ Now, $|\gamma_{k}|\cap f_{m_{k}}(D_{m_{k}})\neq\varnothing\neq|\gamma_{k}|\setminus f_{m_{k}}(D_{m_{k}})$ and by Proposition 3 $|\gamma_{k}|\cap\partial f_{m_{k}}(D_{m_{k}})\neq\varnothing.$ Thus, $h(f_{m_{k}}(x_{0}),\partial f_{m_{k}}(D_{m_{k}}))\rightarrow 0$ as $k\rightarrow\infty,$ that contradicts with $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0},$ proved above. Theorem is proved. $\Box$
6 The versions for prime ends
The following statement may be proved completely similarly to [Sev1, Lemma 5.2], cf. [Ad, Theorem 5, Example 5], [KR, Lemma 5, Theorem 3].
Lemma 6.1.  Let $p\geqslant 1,$ let $D$ be a quasiconformally regular domain, let $f:D\rightarrow\overline{{\mathbb{R}}^{n}}$ be an open discrete and closed ring $Q$-mapping with respect to the $p$-modulus at all points $b\in\partial D,$ let $f(D)=D^{\,\prime}$, and let $D^{\,\prime}$ be a domain with a strongly accessible boundary with a respect to $p$-modulus at least at one point $y\in C(f,b).$ Assume that, for any $b\in\partial D$ there is $\varepsilon_{0}>0$ and some positive measurable function $\psi:(0,\varepsilon_{0})\rightarrow(0,\infty)$ such that
$0<I(\varepsilon,\varepsilon_{0})=\int\limits_{\varepsilon}^{\varepsilon_{0}}\psi(t)\,dt<\infty$
for any $\varepsilon\in(0,\varepsilon_{0})$ and, in addition,
$\int\limits_{A(b,\varepsilon,\varepsilon_{0})}Q(x)\cdot\psi^{\,p}(|x-b|)\ dm(x)=o(I^{p}(\varepsilon,\varepsilon_{0}))\,$,
where $A:=A(b,\varepsilon,\varepsilon_{0})$ is defined in (1.2). Then $f$ have a continuous extension $\overline{f}:\overline{D}_{P}\rightarrow\overline{D^{\,\prime}}$ such that $f(\overline{D}_{P})=\overline{D^{\,\prime}}.$
The following statement holds, cf. [SevSkv].
Lemma 6.2.   The statement of Theorem 1 remains true, if, under assumptions of Theorem 1, instead of the assumption 4) we require the following: there exists $\varepsilon_{0}=\varepsilon_{0}(x_{0})>0$ and a Lebesgue measurable function $\psi:(0,\varepsilon_{0})\rightarrow[0,\infty]$ such that (2.1) holds while the relation (2.2) holds as $\varepsilon\rightarrow 0,$ where $A(x_{0},\varepsilon,\varepsilon_{0})$ is defined in (1.2).
Proof.   I. Assume the contrary. Then there exists $a>0$ such that for $\delta=1/k,$ $k=1,2,\ldots,$ there are $m_{k}\in{\mathbb{N}}$ and $x_{k},x^{\,\prime}_{k}\in D_{m_{k}}$ such that $\rho(x_{k},x^{\,\prime}_{k})<1/k,$ however, $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a.$ By the definition, $\overline{D_{0}}_{P}$ is a compact space, so that we may consider that the sequences $x_{k},x^{\,\prime}_{k}\rightarrow P_{0}\in\overline{D_{0}}_{P}$ by the metrics $\rho.$ We may consider that the sequence $m_{k}$ is increasing by $k.$ Otherwise, $m_{k}=k_{0}$ for sufficiently large $k\in{\mathbb{N}}$ and some $k_{0}\in{\mathbb{N}},$ besides that, $P_{0}\in\overline{D_{k_{0}}}_{P}$ because $f_{k_{0}}$ is defined at $x_{k}\cap D_{k_{0}}$ and $x_{k}\rightarrow P_{0}$ as $k\rightarrow\infty.$ Now, from the assumption above we have that $h(f_{k_{0}}(x_{k}),f_{k_{0}}(x^{\,\prime}_{k}))\geqslant a.$ The latter contradicts with Lemma 6. So, $m_{k}$ may be chosen as an increasing sequence. Without loss of generality, going to renumbering, if required, we may consider that the relation $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a$ holds for $m$ instead of $m_{k},$ i.e., $x_{m},x^{\,\prime}_{m}\in D_{m}$ and
$h(f_{m}(x_{m}),f_{m}(x^{\,\prime}_{m}))\geqslant a/2\qquad\forall\,\,m\in{\mathbb{N}}\,.$ (6.1)
If $P_{0}\in D_{0},$ then all mappings $f_{m}$ except of finite numbers are defined in some ball $B(P_{0},\varepsilon_{0})\subset D_{0}\cap D_{m},$ $m\geqslant M_{0}.$ Now, $f_{m}$ is equicontinuous at $P_{0}$ (see [RS1, Lemma 4.9] for $p=n$ and [Sev1, Lemma 4.3] for n1<p<n.n-1<p<n. Therefore, we may consider that $P_{0}\in E_{D_{0}}=\overline{D_{0}}_{P}\setminus D_{0}.$
Let $d_{m}$ and $\sigma_{m},$ $m=0,1,2,\ldots,$ be sequences of domains and cuts in $P_{0},$ respectively, while $\sigma_{m}$ lie on the spheres $S(z_{0},r_{m})$ centered at some point $z_{0}\in\partial D_{0},$ where $r_{m}\rightarrow 0$ as $m\rightarrow\infty$ (such a sequence $\sigma_{m}$ exists by [IS, Lemma 3.1], cf. [KR]). Since $D_{m}$ is a regular sequence of domains with a respect to $D_{0}$ in the terms of prime ends, there exists a sequence of cuts $\varsigma_{m}$ and domains $d^{\,\prime}_{m},$ equivalent to $\sigma_{m}$ and $d_{m},$ respectively, such that the following condition holds: given $d^{\,\prime}_{k},$ $k=1,2,\ldots,$ there is $M_{1}=M_{1}(k)$ such that $d^{\,\prime}_{k}\cap D_{m}$ is a non-empty connected set for every $m\geqslant M_{1}(k).$ Now, we will consider that $d^{\,\prime}_{k}\subset d_{k}$ for any $k=1,2,\ldots.$
Thus, given a domain $d^{\,\prime}_{1}$ there is a number $M(1)\in{\mathbb{N}}$ such that $D_{m}\cap d^{\,\prime}_{1}$ is connected for $m\geqslant M(1).$ Since $x_{m},x_{m}^{\,\prime}\rightarrow P_{0}$ as $m\rightarrow\infty,$ there is $m_{1}\geqslant M(1)$ such that $x_{m_{1}},x^{\,\prime}_{m_{1}}\in d^{\,\prime}_{1}\cap D_{m_{1}}.$ Follow, given a domain $d^{\,\prime}_{2}$ there is a number $M(2)\in{\mathbb{N}}$ such that $D_{m}\cap d^{\,\prime}_{2}$ is connected for $m\geqslant M(2).$ Since $x_{m},x_{m}^{\,\prime}\rightarrow P_{0}$ as $m\rightarrow\infty,$ there is $m_{2}\geqslant M(2)$ such that $x_{m_{2}},x^{\,\prime}_{m_{2}}\in d^{\,\prime}_{2}\cap D_{m_{2}}.$ Etc. In general, given $k\in{\mathbb{N}}$ and a domain $d^{\,\prime}_{k}$ there is a number $M(k)\in{\mathbb{N}}$ such that $D_{m}\cap d^{\,\prime}_{k}$ is connected for $m\geqslant M(k).$ Since $x_{m},x_{m}^{\,\prime}\rightarrow P_{0}$ as $m\rightarrow\infty,$ there is $m_{k}\geqslant M(k)$ such that $x_{m_{k}},x^{\,\prime}_{m_{k}}\in d^{\,\prime}_{k}\cap D_{m_{k}}.$ Relabeling, if required, we may consider that the same sequences $x_{m}$ and $x^{\,\prime}_{m}$ satisfy the above conditions, i.e., $x_{m},x^{\,\prime}_{m}\in d^{\,\prime}_{m}\cap D_{m}$ while $d^{\,\prime}_{m}\cap D_{m}$ is connected. We join the points $x_{m}$ and $x^{\,\prime}_{m}$ by a path $\gamma_{m}:[0,1]\rightarrow D_{m}\cap d^{\,\prime}_{m}$ such that $\gamma_{m}(0)=x_{m},$ $\gamma_{m}(1)=x^{\,\prime}_{m}$ and $\gamma_{m}(t)\in D_{m}\cap d^{\,\prime}_{m}$ for $t\in(0,1),$ see Figure 3.
Figure 3: To prove Lemma 2
Denote by $C_{m}$ the image of the path $\gamma_{m}(t)$ under $f_{m}.$ It follows from relation (6.1) that
$h(C_{m})\geqslant a/2\qquad\forall\,m\in{\mathbb{N}}\,$, (6.2)
where $h$ denotes the chordal diameter of the set.
Let $\Gamma_{m}$ be a family of paths joining $\gamma_{m}$ and $A$ in $D_{m}.$ Observe that, $A\subset D_{0}\setminus d_{m}$ for sufficiently large $m\in{\mathbb{N}}.$ We may consider that the latter holds for all $m\in{\mathbb{N}}.$ Now, by Proposition 3
$\Gamma_{m}>\Gamma(S(z_{0},r_{m}),S(z_{0},r_{1}),D_{0})\,.$ (6.3)
By the definition of the class $\mathfrak{F}_{Q,A,p,\delta}(D_{0},\mathfrak{D}),$ the relations (1.3)–(1.4) hold at $z_{0}.$ Therefore, by (6.3)
$M_{p}(f_{m}(\Gamma_{m}))\leqslant\int\limits_{A(z_{0},r_{m},r_{1})}Q(x)\cdot\eta^{p}(|x-z_{0}|)\ dm(x)$ (6.4)
for every Lebesgue measurable function $\eta:(r_{m},r_{1})\rightarrow[0,\infty]$ such that $\int\limits_{r_{m}}^{r_{1}}\eta(r)\,dr\geqslant 1.$ Observe that, the function
[…]
where $I(r_{m},r_{1}):=\int\limits_{r_{m}}^{r_{1}}\psi(t)\,dt.$ Therefore, it follows from (2.2) and (6.4) that
$M_{p}(f_{m}(\Gamma_{m}))\leqslant\alpha(r_{m})\rightarrow 0$ (6.5)
as $m\rightarrow\infty,$ where $\alpha(\varepsilon)$ is some nonnegative function tending to zero as $\varepsilon\rightarrow 0,$ which exists due to the condition (2.2).
On the other hand, observe that $f_{m}(\Gamma_{m})=\Gamma(C_{m},f_{m}(A_{m}),D_{m}^{\,\prime}).$ By the hypothesis of the lemma, $h(f_{m}(A_{m}))\geqslant\delta$ for all $m\in{\mathbb{N}}.$ Therefore, due to (6.2), $h(f_{m}(A_{m}))\geqslant\delta_{1}$ and $h(C_{m})\geqslant\delta_{1},$ where $\delta_{1}:=\min\{\delta,a/2\}.$ Using the fact that the domains $D_{m}^{\,\prime}:=f_{m}(D_{m})$ are equi-uniform with respect to the $p$-modulus, we conclude that there exists $\sigma>0$ such that
$M_{p}(f_{m}(\Gamma_{m}))=M_{p}(\Gamma(C_{m},f_{m}(A_{m}),D_{m}^{\,\prime}))\geqslant\sigma\qquad\forall\,\,m\in{\mathbb{N}}\,$,
which contradicts the condition (6.5). The obtained contradiction indicates that the assumption of the absence of equicontinuity of the family $\mathfrak{F}_{Q,A,p,\delta}(D_{0},\mathfrak{D})$ at $z_{0}$ was incorrect. The obtained contradiction completes the proof of the uniform equicontinuity of $f_{m},$ $m=1,2,\ldots,$ in $\mathfrak{D}$ in terms of prime ends.
[…]
Now, we give the version of Lemma 6 for mappings with a branching.
Lemma 6.3.   The statement of Theorem 1 remains true, if, under assumptions of this theorem, instead of the assumption 4) mentioned here we require the following: there exists $\varepsilon_{0}=\varepsilon_{0}(x_{0})>0$ and a Lebesgue measurable function $\psi:(0,\varepsilon_{0})\rightarrow[0,\infty]$ such that (2.1) holds while the relation (2.2) is satisfied as $\varepsilon\rightarrow 0,$ where $A(x_{0},\varepsilon,\varepsilon_{0})$ is defined in (1.2).
Proof.   I. The proof is much the same as that of Lemma 6, so we will limit ourselves to a schematic proof.
Assume the contrary. Then there exists $a>0$ such that for $\delta=1/k,$ $k=1,2,\ldots,$ there are $m_{k}\in{\mathbb{N}}$ and $x_{k},x^{\,\prime}_{k}\in D_{m_{k}}$ such that $\rho(x_{k},x^{\,\prime}_{k})<1/k,$ however, $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a.$ By the definition, $\overline{D_{0}}_{P}$ is a compact space, so that we may consider that the sequences $x_{k},x^{\,\prime}_{k}\rightarrow P_{0}\in\overline{D_{0}}_{P}$ by the metrics $\rho.$ Arguing similarly to the proof of Lemma 6, we may consider that (6.1) holds. If $P_{0}\in D_{0}$, then $f_{m}$ is equicontinuous at $P_{0}$ (see [RS1, Lemma 4.9] for $p=n$ and [Sev1, Lemma 4.3] for n1<p<n.n-1<p<n. Now, we may consider that $P_{0}\in E_{D_{0}}=\overline{D_{0}}_{P}\setminus D_{0}.$ Under notations of Lemma 6 we may consider that $x_{m},x^{\,\prime}_{m}\in d^{\,\prime}_{m}\cap D_{m}$ while $d^{\,\prime}_{m}\cap D_{m}$ is connected. We join the points $x_{m}$ and $x^{\,\prime}_{m}$ by a path $\gamma_{m}:[0,1]\rightarrow D_{m}\cap d^{\,\prime}_{m}$ such that $\gamma_{m}(0)=x_{m},$ $\gamma_{m}(1)=x^{\,\prime}_{m}$ and $\gamma_{m}(t)\in D_{m}\cap d^{\,\prime}_{m}$ for $t\in(0,1).$ Set $C_{m}:=|f_{m}(\gamma_{m})|.$ Now, $h(C_{m})\geqslant a/2$ for sufficiently large $m\in{\mathbb{N}}.$
By the definition of $\mathfrak{R}_{Q,\delta,p,E}(D_{0},\mathfrak{D}),$ for any $f_{m}\in\mathfrak{R}_{Q,\delta,p,E}(D_{0},\mathfrak{D})$ and any domain $D^{\,\prime}_{m}:=f_{m}(D_{m})$ there exists a continuum $K_{m}\subset D^{\,\prime}_{m}$ such that $h(K_{m})\geqslant\delta$ and $h(f_{m}^{\,-1}(K_{m}),\partial D_{m})\geqslant\delta>0.$ Since by the assumption the domains $D^{\,\prime}_{m}$ form the equi-uniform family with respect to $p$-modulus and $h(C_{m})\geqslant a/2$ we obtain that
$M_{p}(\Gamma(K_{m},C_{m},D^{\,\prime}_{m}))\geqslant b\,$, (6.6)
$m=1,2,\ldots,$ for some $b>0.$ Consider the family $\Gamma_{m},$ consisting of all paths $\beta:[0,1)\rightarrow D^{\,\prime}_{m},$ where $\beta(0)\in C_{m}$ and $\beta(t)\rightarrow p\in K_{m}$ as $t\rightarrow 1.$ Let $\Gamma^{*}_{m}$ be a family of all total $f_{m}$-liftings $\alpha:[0,1)\rightarrow D$ of the family $\Gamma_{m}$ with origin at $|\gamma_{m}|$ (see Proposition 3). Since $f_{m}$ is closed, $\alpha(t)\rightarrow f^{\,-1}_{m}(K_{m}),$ where $f^{\,-1}_{m}(K_{m})$ is the total preimage of the continuum $K_{m}$ under the mapping $f_{m}.$ Since $h(f_{m}^{\,-1}(K_{m}),\partial D_{m})\geqslant\delta>0$ and $D_{m}\subset D_{0},$ by Proposition 4, $h(f_{m}^{\,-1}(K_{m}),\partial D_{0})\geqslant\delta>0,$ as well. Now, we may consider that $f_{m}^{\,-1}(K_{m})\subset D_{0}\setminus d^{\,\prime}_{1}.$ In this case, by Proposition 3
$\Gamma^{*}_{m}>\Gamma(S(z_{0},r_{m}),S(z_{0},r_{1}),D_{0})\,.$ (6.7)
It follows by (6.7) and (1.3) that,
$M_{p}(f_{m}(\Gamma_{m}^{*}))\leqslant M_{p}(\Gamma(S(z_{0},r_{m}),S(z_{0},r_{1}),D_{0}))$
$\leqslant M_{p}(f_{m}(\Gamma(|\gamma_{m}|,f^{\,-1}_{m}(K_{m}),D_{m})))\leqslant\int\limits_{A(x_{0},r_{m},r_{1})}Q(x)\cdot\eta^{p}(|x-z_{0}|)\ dm(x)$ (6.8)
for each measurable function $\eta:(r_{m},r_{1})\rightarrow[0,\infty],$ such that $\int\limits_{r_{1}}^{r_{m}}\eta(r)\,dr\geqslant 1.$ Again, arguing similarly to the proof of Lemma 6, we obtain from (6.8) that
$M_{p}(f_{m}(\Gamma^{\,*}_{m}))\leqslant\alpha(r_{m})\rightarrow 0$ (6.9)
as $m\rightarrow\infty,$ where $\alpha(\varepsilon)$ is some nonnegative function tending to zero as $\varepsilon\rightarrow 0$ (which exists by the condition (2.2)). In addition, $f_{m}(\Gamma^{\,*}_{m})=\Gamma_{m}$ and $M_{p}(\Gamma_{m})=M_{p}(\Gamma(K_{m},C_{m},D^{\,\prime}_{m})).$ Thus
$M_{p}(f_{m}(\Gamma^{\,*}_{m}))=M_{p}(\Gamma(K_{m},C_{m},D^{\,\prime}_{m}))\,.$ (6.10)
However, relations (6.9) and (6.10) contradict with (6.6). The resulting contradiction indicates that the assumption (6.1) was incorrect, and, therefore, the family $\mathfrak{R}_{Q,\delta,p,E}(D_{0},\mathfrak{D})$ is equicontinuous at $P_{0}\in E_{D_{0}}.$
[…]
Example 1. Let $D_{0}$ be the unit disk $\mathbb{D}$} in $\mathbb{C}$} with a cut along the segment $[0,1]$, i.e., $D_{0}={\mathbb{D}}\setminus I\subset{\mathbb{C}}$, where ${\mathbb{D}}:=\{z\in{\mathbb{C}}\,:\,|z|<1\}$, $I:=\{z=x+iy\in{\mathbb{C}}\,:\,y=0,0\leqslant x\leqslant 1\}$. Let $f_{m}$ be a conformal mapping of $D_{m}=\{z\in{\mathbb{C}}\,:\,|z|<m/(m+1)\}\setminus I_{m},$ $I_{m}:=\{z=x+iy\in{\mathbb{C}}\,:\,y=0,0\leqslant x\leqslant m/(m+1)\},$ onto $\mathbb{D}$}. For instance, we may take
$f_{m}(z)=\frac{\left(\frac{\sqrt{zr_{m}}+1}{\sqrt{zr_{m}}-1}\right)^{2}-i}{\left(\frac{\sqrt{zr_{m}}+1}{\sqrt{zr_{m}}-1}\right)^{2}-i}\,,\quad z=x+iy\,,\quad i^{2}=-1\,$,
where $\sqrt{z}$ denotes the main branch of the square root and $r_{m}=(m+1)/m.$ We verify conditions of Theorem 1. First of all, $f_{m}$ are homeomorphisms of $D_{m}$ onto $\mathbb{D}$} satisfying (1.3)–(1.4) at any point of $\overline{D_{0}}=\overline{\mathbb{D}}$ with $Q\equiv 1$ and $p=n=2$ as conformal mappings (see [Va, Theorem 8.1]). Since this sequence (obviously) converges to the conformal mapping $f(z)=\frac{\left(\frac{\sqrt{z}+1}{\sqrt{z}-1}\right)^{2}-i}{\left(\frac{\sqrt{z}+1}{\sqrt{z}-1}\right)^{2}-i}$ locally uniformly in $D_{0},$ the relation $h(f_{m}(A))\geqslant\delta$ obviously holds for any $m=1,2,\ldots,$ any continuum $A$ in $D_{0}\cap\bigcap\limits_{m=1}^{\infty}D_{m}$ and some $\delta>0.$ Besides that, the relation $h(\overline{{\mathbb{R}}^{n}}\setminus f_{m}(D_{m}))=h(\overline{{\mathbb{R}}^{n}}\setminus{\mathbb{D}})\geqslant\delta$ holds, for in0stance, with $\delta=1.$ Observe that $D_{m},$ $m=1,2,\ldots,$ converge to $D_{0}={\mathbb{D}}\setminus I$ as its kernel, see Figure 4.
Figure 4: Illustration for Example 6.
Besides that, the domains $d_{k},$ formed by chains of cuts lying on the circles centered at some point $z_{0}\in\partial{\mathbb{D}}\cup I$ form a connected intersection with the domains $D_{m}$ for each fixed $k$ and all sufficiently large $m.$ Indeed, prime ends of $D_{0}$ may be conditionally divided into 6 groups: 0) the set of prime ends $P_{0}$ which correspond to $z_{0}\in{\mathbb{S}}^{1}\setminus I,$ where ${\mathbb{S}}^{1}=\partial{\mathbb{D}};$ 1) and 2): the prime ends $P_{1}$ and $P_{2}$ with impression $z_{0}=1,$ the sequences of domains $d_{k}$ lie from below and from upper of $I,$ correspondingly; 3) and 4): the set of prime ends $P_{3}$ and $P_{4}$ with impression at some $z_{0}\in I\setminus\{1,0\},$ the sequences of domains $d_{k}$ lie from below and from upper of $I,$ correspondingly; 5) the prime end $P_{5}$ which correspond to its impression $I(P_{5})=\{0\}.$ Let $d_{k},$ $k=1,2,\ldots,$ be a sequence of domains in $P_{i},$ $i=\overline{0,5}.$ Now, $d_{k}$ is convex for any $k=1,2,\ldots.$ Given $k\in{\mathbb{N}},$ the intersection $D_{m}\cap d_{k}$ is non-empty and convex for sufficiently large $m.$ In particular, $D_{m}\cap d_{k}$ is connected for above $m,$ as required. Thus, 1) the sequence of domains $D_{m}$ is regular with respect to $D_{0}$ in terms of prime ends. Observe that, 2) for every $m\in{\mathbb{N}},$ a domain $D_{m}$ is quasiconformally regular; $D_{0}$ is also quasiconformally regular. Indeed, $D_{0}$ and $D_{m}$ are regular because by the Riemannian mapping theorem, $D_{0}$ (or $D_{m}$) is conformally equivalent to $\mathbb{D}$}; moreover, $\mathbb{D}$} has locally quasiconformal boundary (see e.g. [Va, Theorem 17.10]). Note that, 3) the family $f_{m}(D_{m})={\mathbb{D}}$ is equi-uniform with respect to $2$-modulus over all $m\in{\mathbb{N}}$ (see Remark 1). Finally, 4) $Q\equiv 1$ has a finite mean oscillation in $\overline{D_{0}};$ moreover, (1.5) holds for every $x_{0}\in\overline{D_{0}}$ and any $\beta(x_{0})>0.$ Thus, all of conditions of Theorem 1 hold, as required.
Example 2. It is possible to construct corresponding families of mappings, satisfying Theorem 1 that have unbounded characteristics. Let $z_{0}\in D_{0}$ and $0<r_{0}<d(z_{0},\partial D_{0}),$ where $D_{0}$ is a domain from Example 6. Put $h(x)=\frac{x}{|x|\log\frac{(r_{0}e)}{|x|}},$ $x\in B(z_{0},r_{0}),$ $h(z_{0})=z_{0},$ $h|_{S(z_{0},r_{0})}=x.$ Then $h$ is defined in the ball $B(z_{0},r_{0})$ and $h(B(z_{0},r_{0}))=B(z_{0},r_{0}).$ Reasoning similarly to [MRSY, Proposition 6.3], it may be shown that $h$ satisfies the relations (1.3)–(1.4) at any point $x_{0}\in\overline{B(z_{0},r_{0})}$ for $Q=Q(x)=\log\left(\frac{r_{0}e}{|x|}\right).$ Note that $Q$ satisfies (1.5) for every $x_{0}\in\overline{D_{0}}$ and any $\beta(x_{0})>0$ that may be verified directly. Now, we put
$g_{m}(x)=\begin{cases}(f_{m}\circ h)(x)\,,&x\in B(z_{0},r_{0})\,,\\ f_{m}(x)\,,&x\not\in B(z_{0},r_{0})\end{cases}\,$,
where $f_{m}$ and $f$ are from Example 6. By the construction, $g_{m}$ satisfy the relations (1.3)–(1.4) at any point $x_{0}\in\overline{D_{0}}$ for $Q=Q(y)=\log\left(\frac{r_{0}e}{|x|}\right).$ Observe that, the domains $f_{m}(D_{m})$ are not changed, so that the domains $g_{m}(D_{m})$ are regular, as well. In addition, since $h$ is a fixed mapping, the relations $h(g_{m}(A))\geqslant\delta,$ $m=1,2,\ldots,$ also hold for some $\delta>0$ and all $m=1,2,\ldots.$ The mappings $g_{m},$ $m=1,2,\ldots,$ are homeomorphisms and satisfy all the conditions of Theorem 1.
Declarations.
Conflict of interest. The authors have no financial or proprietary interests in any material discussed in this article.
Funding. The work was supported by the National Research Foundation of Ukraine, Project number 2025.07/0014, Project name: ‘‘Modern problems of Mathematical Analysis and Geometric Function Theory’’.
Data availability statement. All necessary data are included in the paper.
Contribution. The authors have equal contribution to writing the article.